### v1

# FALCON: Fine-grained Activation Manipulation by Contrastive Orthogonal Unalignment for Large Language Model

Abstract
Large language models have been widely applied, but can inadvertently encode sensitive or harmful information, raising significant safety concerns. Machine unlearning has emerged to alleviate this concern; however, existing training-time unlearning approaches, relying on coarse-grained loss combinations, have limitations in precisely separating knowledge and balancing removal effectiveness with model utility. In contrast, we propose Fine-grained Activation manipuLation by Contrastive Orthogonal uNalignment (FALCON), a novel representation-guided unlearning approach that leverages information-theoretic guidance for efficient parameter selection, employs contrastive mechanisms to enhance representation separation, and projects conflict gradients onto orthogonal subspaces to resolve conflicts between forgetting and retention objectives. Extensive experiments demonstrate that FALCON achieves superior unlearning effectiveness while maintaining model utility, exhibiting robust resistance against knowledge recovery attempts.
Machine Learning, ICML
1 Introduction
Recent advancements in generative AI (Achiam et al., 2023; Anil et al., 2023; Dubey et al., 2024), fueled by innovations in training techniques such as Parameter-Efficient Fine-Tuning (PEFT), have enabled large language models (LLMs) to efficiently internalize linguistic knowledge and excel in tasks from text generation to decision making (Hu et al., 2022; yang Liu et al., 2024). These models derive their power from massive, diverse corpora, but this dependency on large-scale datasets introduces significant risks: harmful, biased, or sensitive information can be inadvertently encoded and amplified, leading to ethical violations, regulatory noncompliance, and potential misuse (Hsu et al., 2024; Urman & Makhortykh, 2023; Jiao et al., 2024).
Existing mitigation strategies, such as implementing guardrails (Dong et al., 2024) or training models with expertly curated datasets to refuse harmful queries (Ouyang et al., 2022), are computationally expensive and often inadequate against adversarial attacks (Xu et al., 2024; Yin et al., 2024). In contrast, while retraining an entire model on a cleaned dataset to eliminate harmful influences is theoretically feasible, it is prohibitively resource-intensive for modern LLMs (Jin et al., 2023). Additionally, adversaries can exploit PEFT to reintroduce such unwanted information, highlighting the urgent need for more effective and scalable solutions for publicly accessed LLMs (Qi et al., 2024).
To solve harmful or sensitive information in machine learning models, Machine Unlearning (MU) has emerged as a promising solution, supported by growing regulations such as the “right to be forgotten” under the GDPR (Regulation, 2016; Ginart et al., 2019). It originally developed in the non-LLMs domain and has proven effective at removing specific data influences while preserving model performance (Bourtoule et al., 2021; Kim & Woo, 2022a; Liu et al., 2023). When transferred to maintain responsible LLMs, MU offers significant advantages, being far more computationally efficient than full retraining. Unlearned models also exhibit greater inherent safety, as they lack the undesired knowledge necessary for malicious behaviors (Hendrycks et al., 2021b; Li et al., 2024).
Despite its potential, LLM unlearning still faces several fundamental issues: (1) existing approaches typically rely on empirical methods like grid search to identify intervention parameters, lacking an efficient and interpretable way to guide parameter selection within deep architectures, (2) current methods normally rely on coarse-grained manipulation (using simplistic loss combinations that induce random representation dispersion with uncontrolled gradient dynamics, struggling to balance knowledge removal and utility preservation) rather than fine-grained representation manipulation (achieving more effective knowledge separation through targeted representation modification and regulated gradient dynamics for reducing damage to model utility), and (3) knowledge recovery methods like jailbreaking attack can recover the undesired information from the unlearned model (Shaik et al., 2024).
To address the aforementioned issues of selective knowledge unlearning in LLMs, we propose Fine-grained Activation manipuLation by Contrastive Orthogonal uNalignment (FALCON), a representation-guided framework designed for precise knowledge removal while maintaining critical retention. FALCON leverages mutual information (MI) to evaluate dependencies between forget and retain data, enabling interpretable parameter selection and optimization guidance. For multi-domain knowledge unlearning, FALCON achieves fine-grained representation unalignment through two mechanisms: directional contrastive unalignment with singular value decomposition (SVD)-derived principal offset vectors and gradient orthogonal projection for precise control of knowledge separation granularity. Furthermore, our framework does not require full access to training data, making it practical for scenarios where such access is limited. FALCON further employs a second-order optimizer with approximate curvature information and can achieve efficient unlearning through single-layer parameter modification, ensuring both computational efficiency and scalability across diverse knowledge domains. Our contributions are as follows:
  • We propose FALCON, a representation-guided framework that combines contrastive mechanisms and gradient projection to achieve fine-grained representation unalignment in LLMs.
  • We introduce information-theoretic metrics for quantifying knowledge entanglement, enabling principled parameter selection and providing theoretical insights into knowledge distribution across model architectures.
  • We demonstrate the scalability, effectiveness, and resistance to knowledge recovery of FALCON through extensive experiments, highlighting its ability to unlearn selective knowledge while preserving utility across various LLMs.
2 Related work
Our paper focuses on LLM unlearning for undesired knowledge, information-theoretic metrics, and contrastive learning. We highlight the developments and limitations of LLM unlearning in this section, while related advancements in information-theoretic metrics, contrastive learning, and gradient projection are detailed in the Appendix A and B.
LLM Unlearning
LLM unlearning refers to the selective removal of specific knowledge from large language models while preserving their overall functionality (Zhang et al., 2024a). Current approaches can be broadly categorized into training-time methods and inference-time methods (Barez et al., 2025). Among training-time approaches, which represent the mainstream methodology, two primary directions have emerged. The first direction focuses on gradient optimization, maximizing the loss function to suppress harmful knowledge through techniques such as gradient ascent (Yao et al., 2023; Jang et al., 2023) and reverse gradients (Eldan & Russinovich, 2023). The second direction emphasizes representation-guided adaptation, targeting specific intermediate hidden representations for modification (Li et al., 2024). While these aforementioned training-time methods achieve permanent unlearning by targeting specific layers and parameters, they currently rely heavily on coarse-grained loss combinations that struggle to disentangle deeply embedded knowledge representations flexibly.
Inference-time methods offer alternative approaches through task vectors and model editing. Task vector approaches address efficiency concerns through arithmetic operations on parameter-efficient modules, enabling lightweight unlearning under resource constraints (Ilharco et al., 2023; Zhang et al., 2023). Model editing usually modifies intermediate hidden states or logits to alter model behavior (Barez et al., 2025), such as contrastive decoding methods that suppress undesired content generation (Zhong et al., 2024). However, these methods’ dependence on modular arithmetic operations fundamentally limits their granularity in knowledge separation and constrains generalizability across diverse scenarios. Additionally, in-context unlearning has emerged as another inference-time approach, leveraging tailored prompts to dynamically suppress undesired outputs (Zheng et al., 2023; Pawelczyk et al., 2024). While flexible, this method’s effect remains inherently temporary as the undesired knowledge persists in the model’s representation space (Liu et al., 2024b).
Despite these advancements, existing training-time methods fall short in achieving precise knowledge disentanglement between information to be forgotten and retained. To address these limitations, we propose FALCON, a targeted representation unalignment approach that achieves more precise separation through contrastive learning, gradient projection, and information-theoretic guidance. Through its contrastive mechanism and gradient projection, our approach enables fine-grained knowledge separation and resolves optimization conflicts between forgetting and retention objectives, while enhanced resistance compared to current state-of-the-art training-time methods.
3 Problem Formulation
3.1 Problem Setup
Figure 1: Schematic representation of the proposed FALCON method. The pipeline involves parameter selection based on mutual information, the construction of principal offset vectors, the application of contrastive mechanism on both datasets, and gradient conflict resolution to balance forgetting and retaining objectives.
The task of LLM unlearning involves selectively removing specific knowledge (forget set) from the model while retaining critical information (retain set). However, this process is complicated by the issue of knowledge entanglement, where representations of the forget and retain sets overlap significantly within the model’s parameters (Zhang et al., 2024b). This entanglement arises due to the distributed nature of knowledge across multiple layers and features, making it difficult to isolate knowledge for removal without affecting retained information.
To formalize the unlearning process, we adopt the general formulation proposed by Liu et al. (Liu et al., 2024b):
$\min_{\theta}\left\{\mathbb{E}_{(x,y_{f})\in\mathcal{D}_{\mathcal{F}}}\left[\mathcal{L}(y_{f}|x;\theta)\right]+\lambda\mathbb{E}_{(x,y)\in\mathcal{D}_{\mathcal{R}}}\left[\mathcal{L}(y|x;\theta)\right]\right\}$ (1)
where $\mathcal{L}(y|x;\theta)$ measures the discrepancy between the model’s prediction and the target response $y$ for a given input $x$ under the model’s parameters $\theta$. Here, $\mathcal{D}_{\mathcal{F}}$ and $\mathcal{D}_{\mathcal{R}}$ denote the forget set and retain set, respectively. The variable $y_{f}$ specifies the intended output for the forget set after unlearning, while the hyperparameter $\lambda\geq 0$ controls the trade-off between forgetting and retention objectives. For simplicity, we will refer to this objective as $\min_{\theta}\mathbb{E}_{\text{MU}}(\theta)$ in subsequent sections.
Despite the generality of above formulation, it does not explicitly quantify the representations of forgotten and retained knowledge. This lack of quantification poses challenges in precisely guiding the unlearning process (Qu et al., 2024). To address this, a principled metric is needed to evaluate and minimize knowledge entanglement, ensuring that unlearning primarily affects the forget set while minimizing interference with the retain set.
To address this, we introduce information-theoretic measures, specifically continuous entropy and mutual information, to quantify the dependency between the activations of the forget and retain sets. Let $\mathcal{F}$ and $\mathcal{R}$ represent the activations of the forget and retain sets at a specific layer of the model, respectively. The degree of knowledge entanglement between representations can be formulated as the mutual information $I(\mathcal{F};\mathcal{R})$:
$I(\mathcal{F};\mathcal{R})=H(\mathcal{F})+H(\mathcal{R})-H(\mathcal{F},\mathcal{R})$ (2)
where $H(\mathcal{F})$ and $H(\mathcal{R})$ are the continuous entropies of the activations $\mathcal{F}$ and $\mathcal{R}$, and $H(\mathcal{F},\mathcal{R})$ denotes their joint entropy. These measures provide a systematic approach to identify layers with minimal entanglement and guiding the optimization process for targeted and effective unlearning. The details of these metrics are shown in Appendix C.
3.2 LLM unlearning with MI Guidance
To quantify knowledge entanglement during machine unlearning, we use MI to measure the dependency between the activations of the forget set $\mathcal{F}^{(l)}$ and the retain set $\mathcal{R}^{(l)}$ at each layer $l$. The MI $I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$ serves as an indicator to guide the unlearning process by minimizing entanglement between $\mathcal{F}^{(l)}$ and $\mathcal{R}^{(l)}$.
To minimize the entanglement between the forget and retain sets’ representations, we formulate the layer selection as:
$l^{*}=\arg\min_{l}I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$ (3)
Given the selected layer $l^{*}$, the LLM unlearning problem guided by MI can be reformulated as:
[…]
This formulation ensures that the unlearning process is conducted on the parameters with minimal knowledge entanglement, effectively suppressing the undesired knowledge while reducing interference with the retained knowledge.
4 Methodology
To address the challenges of more thorough selective multi-domain knowledge unlearning and enhanced robustness against knowledge recovery in LLMs, we propose FALCON shown in Figure 1, a framework that advances both precision and effective in knowledge manipulation. Unlike prior approaches that rely on coarse-grained loss combinations, FALCON introduces three key mechanisms: (1) mutual information-based guidance to identify parameters where knowledge representations are least entangled, enabling interpretable parameter selection; (2) contrastive mechanism with enhanced representation separation to achieve fine-grained knowledge manipulation while ensuring robust resistance against knowledge recovery attempts; and (3) gradient orthogonal projection to resolve optimization conflicts and ensure training stability. This holistic design enables precise, interpretable, and robust knowledge unlearning in LLMs, transcending traditional loss-combination methods.
4.1 Information-Theoretic Guidance for Unlearning
In this paper, we utilize a principled approach to selective multi-domain knowledge unlearning in LLMs through mutual information. MI provides a natural measure of representational entanglement between the forget and retain datasets across model layers. By identifying layers that minimize MI, we can target unlearning interventions where forget and retain representations exhibit minimal overlap, thus preserving desired knowledge while selectively removing unwanted information.
We extend this measure to the multi-domain scenario where the forget set $\mathcal{F}$ consists of multiple sub-domains $\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{m}$. Our approach quantifies two critical relationships: (1) the interaction between each sub-domain and the retain set $\mathcal{R}$, measured by $I(\mathcal{F}_{i}^{(l)};\mathcal{R}^{(l)})$ at layer $l$, where lower values indicate reduced entanglement and thus more selective unlearning; and (2) the inter-domain dependencies captured by $I(\mathcal{F}_{i}^{(l)};\mathcal{F}_{j}^{(l)})$ for sub-domains $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ $(i\neq j)$, which characterizes potential conflicts or redundancies that may impact unlearning effectiveness.
To quantify the overall representational conflicts between the forget and retain datasets, $I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$, and the interdependence among forgettable sub-domains, $I(\mathcal{F}_{i}^{(l)};\mathcal{F}_{j}^{(l)})$ at layer $l$, we define the aggregate MI as $I^{(l)}$:
$I^{(l)}=\sum_{i=1}^{m}I(\mathcal{F}i^{(l)};\mathcal{R}^{(l)})+\eta\sum_{i=1}^{m}\sum_{j=i+1}^{m}I(\mathcal{F}_{i}^{(l)};\mathcal{F}_{j}^{(l)})$ (5)
where $m$ denotes the number of sub-domains in the forget set $\mathcal{F}$, and $\eta$ is a balancing coefficient that controls the relative importance of inter-domain dependencies.
For each layer $l$, since the activations are high-dimensional and continuous, direct entropy calculation is infeasible (Tsur et al., 2024). Instead, we utilize Kernel Density Estimation (KDE) to approximate the underlying global data distribution, estimating continuous entropy in activation space as defined in Appendix C (Walters-Williams & Li, 2009).
We implement KDE using a Multivariate Gaussian Kernel, which provides a smooth density estimation suitable for high-dimensional data. The estimated probability density function for activations $\mathcal{A}$ is given by:
$p(a)=\frac{1}{Nh}\sum_{n=1}^{N}K\left(\frac{a-a_{n}}{h}\right)$ (6)
where $a\in\mathbb{R}^{d}$ represents a single sample from the activations $\mathcal{A}$, including $\mathcal{F}$ and $\mathcal{R}$, with $d$ denoting the feature dimensionality of the activations, $N$ as the number of samples, $K(\cdot)$ represents the kernel function and $h$ as the adaptive bandwidth calculated using Scott’s rule (Scott, 2015), is defined as $h=\sigma N^{-\frac{1}{d+4}}$, which is particularly suitable for high-dimensional data due to its adjustment based on dimensionality. In this formula, $\sigma$ is the standard deviation of the data. This adaptive bandwidth selection can effectively balance bias and variance, ensuring robust density estimation for diverse activation distributions (Belhaj, 2024). Furthermore, to mitigate the curse of dimensionality, we apply Principal Component Analysis (PCA) to reduce the dimensions of the activations before performing KDE (Altman & Krzywinski, 2018). The number of components is chosen to retain at least 95% of the variance in the activation data, ensuring minimal information loss while significantly lowering computational complexity.
Using the KDE-based entropy estimations, we approximate the overall mutual information $\tilde{I}$ at each layer based on Eq. (5). The optimal layer $l^{*}$ for unlearning is then determined by minimizing $\tilde{I}$:
$l^{*}=\arg\min_{l}\tilde{I}^{(l)}$ (7)
By identifying the layer with the lowest MI, we locate the model region where the forget and retain datasets are least entangled, minimizing the overlap between the two types of knowledge. Concurrently, this layer exhibits higher entanglement among sub-domains within the forget set, enabling efficient updates to shared representations across forgettable sub-domains. This dual property makes the layer the optimal focus for unlearning. During this process, its parameters are selectively optimized to remove undesired knowledge while preserving critical information, ensuring both precision in targeting the forget set and in retaining essential knowledge for downstream tasks.
4.2 Contrastive Orthogonal Unalignment
To achieve selective knowledge unlearning in LLMs, we devise Contrastive Orthogonal Unalignment through contrastive mechanisms and gradient projection to balance knowledge unlearning and retention.
4.2.1 Contrastive Representation Unlearning
The core task of LLM unlearning is to selectively separate knowledge representations to be forgotten from those to be retained. Contrastive learning provides an effective mechanism for this task by learning discriminative representations through comparing similar and dissimilar samples. In our context, we leverage contrastive learning to maximize the distance between representations that should be forgotten while maintaining the coherence of retained knowledge.
To enhance unlearning, we construct Principal Offset Vectors (POV), which steer model activations away from undesired knowledge by reducing their alignment with dominant components identified via SVD in the representation space. The goal of POVs is guiding activations into less entangled subspaces, effectively separating the forget and retain sets.
Mathematically, given an activation matrix $\mathcal{H}\in\mathbb{R}^{(B\cdot L)\times D}$, where $B$ is the batch size, $L$ the sequence length, and $D$ the hidden dimension, we perform SVD to obtain the dominant principal directions ${v_{1},\dots,v_{K}}$ corresponding to the top-$K$ singular values. The POVs $\mathcal{H}^{+}$ is defined as:
$\mathcal{H}^{+}=\frac{f\left(r\cdot\left(I-w\sum_{i=1}^{K}v_{i}v_{i}^{\top}\right),\epsilon\right)}{\left|f\left(r\cdot\left(I-w\sum_{i=1}^{K}v_{i}v_{i}^{\top}\right),\epsilon\right)\right|}$ (8)
Here, $r\in\mathbb{R}^{D}$ is a vector initialized with random values, $w$ controls the influence of principal directions, and $I\in\mathbb{R}^{D\times D}$ is the identity matrix. The term $\epsilon$ introduces optional perturbations. The function $f(\cdot)$ represents a generalized transformation operator like projection and nonlinear mapping, designed to enhance disentanglement and make representations more resistant to recovery.
This formulation ensures $\mathcal{H}^{+}$ is directed away from dominant principal subspaces, combining deterministic guidance and randomized transformations to improve robustness. Unlike generic random vectors, POVs target multiple dominant features, enhancing resistance to adversarial prompts and improving unlearning effectiveness.
For each input sample, we define three types of representations: the anchor representation $\mathcal{H}_{a}$ from the updated model for the forget set, the positive representation $\mathcal{H}^{+}$, given by the POV defined in Equation 8, and the negative representations $\mathcal{H}^{-}$ from the frozen model. To ensure consistent scaling, all representations are normalized, and their similarity scores are measured using cosine similarity:
$S^{+}$ $=\sum_{d=1}^{D}\mathcal{H}_{a}[d]\cdot\mathcal{H}^{+}[d]$
$S^{-}$ $=\sum_{z=1}^{\mathcal{Z}}\sum_{d=1}^{D}\mathcal{H}_{a}[d]\cdot\mathcal{H}_{z}^{-}[d]$ (9)
where $\mathcal{Z}$ is the number of negative samples. Building on these similarity scores, we define the forget loss $\mathcal{L}_{\mathcal{F}}$ using the InfoNCE objective:
$\mathcal{L}_{\mathcal{F}}=-\frac{1}{|B|}\sum_{b=1}^{|B|}\log\frac{\exp(S_{b}^{+}/\tau)}{\exp(S_{b}^{+}/\tau)+\sum_{b=1}^{\mathcal{N}}\exp(S_{b}^{-}/\tau)}$ (10)
where $\tau$ is a temperature scaling parameter. This loss encourages the updated model’s representations to align with the Principal Offset Vectors while diverging from the frozen model’s representations of undesired knowledge. By leveraging both directional guidance through Principal Offset Vectors and contrastive learning, our approach achieves more precise and efficient representation unalignment in activation space.
In addition to unlearning undesired representations, it is essential to preserve critical knowledge required for downstream tasks. To achieve this, we define a retain loss that measures the alignment between the updated model’s activations ($\mathcal{H}^{u}$) and the frozen model’s activations ($\mathcal{H}^{f}$) for the retain set. This alignment is evaluated using a retention alignment loss, which can be viewed as a self-supervised variant of contrastive loss. By maximizing the alignment consistency between the updated and frozen model activations, this loss ensures that critical knowledge is retained effectively during the unlearning process. The retain loss $\mathcal{L}_{\mathcal{R}}$ is defined as:
$\small\mathcal{L}_{\mathcal{R}}=1-\frac{1}{|B|}\sum_{b=1}^{|B|}\frac{\sum_{d=1}^{D}\mathcal{H}_{b}^{u}[d]\cdot\mathcal{H}_{b}^{f}[d]}{\sqrt{\sum_{d=1}^{D}\left(\mathcal{H}_{b}^{u}[d]\right)^{2}}\cdot\sqrt{\sum_{d=1}^{D}\left(\mathcal{H}_{b}^{f}[d]\right)^{2}}}$ (11)
This loss ensures alignment between the updated and frozen model activations for the retain set, preserving critical knowledge while complementing the unlearning objective. Combined with the forget loss $\mathcal{L}_{\mathcal{F}}$, this approach achieves an effective balance between unlearning and retention.
4.2.2 Orthogonalizing Gradient Conflict
After computing the forget loss $\mathcal{L}_{\mathcal{F}}$ and retain loss $\mathcal{L}_{\mathcal{R}}$, we address the misalignment between the optimization directions of unlearning and retaining by employing a gradient projection mechanism. This mechanism orthogonalizes conflicting gradients onto a subspace, minimizing interference and promoting balanced optimization.
Given the gradients of the forget and retain losses, denoted as $\nabla\mathcal{L}_{\mathcal{F}}$ and $\nabla\mathcal{L}_{\mathcal{R}}$, respectively, the conflict can be quantified using the cosine similarity:
$cos(\nabla\mathcal{L}_{\mathcal{F}},\nabla\mathcal{L}_{\mathcal{R}})=\frac{\nabla\mathcal{L}_{\mathcal{F}}\cdot\nabla\mathcal{L}_{\mathcal{R}}}{\|\nabla\mathcal{L}_{\mathcal{F}}\|\cdot\|\nabla\mathcal{L}_{\mathcal{R}}\|}$ (12)
where $cos(\cdot)<0$ indicates opposing directions, signifying a conflict between the two objectives. To mitigate this conflict, we adjust the gradients by projecting one onto the orthogonal complement of the other. Specifically, if $cos(\cdot)<0$, we project $\nabla\mathcal{L}_{\mathcal{F}}$ onto the subspace orthogonal to $\nabla\mathcal{L}_{\mathcal{R}}$:
$\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}=\nabla\mathcal{L}_{\mathcal{F}}-\frac{\nabla\mathcal{L}_{\mathcal{F}}\cdot\nabla\mathcal{L}_{\mathcal{R}}}{\|\nabla\mathcal{L}_{\mathcal{R}}\|^{2}}\nabla\mathcal{L}_{\mathcal{R}}$ (13)
This adjustment ensures that $\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}$ is orthogonal to $\nabla\mathcal{L}_{\mathcal{R}}$, eliminating interference from the retain objective during the update for the forget objective. Once the gradients are adjusted, the final update direction of the FALCON is determined by combined gradients:
$\nabla\mathcal{L}_{FALCON}=\alpha\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}+\beta\nabla\mathcal{L}_{\mathcal{R}}$ (14)
where $\alpha$ and $\beta$ are hyperparameters balancing the contributions of the forget and retain objectives.
This mechanism mitigates gradient conflicts, enabling joint optimization while minimizing interference. By enforcing orthogonality between adjusted gradients, it approximates a Pareto-optimal solution. The model then updates its weights using the conflict-reduced gradient, allowing for more flexible adaptation. To further enhance efficiency and stability, we leverage the second-order optimizer Sophia (Liu et al., 2024a), as suggested in (Gu et al., 2024; Jia et al., 2024), for refined weight updates, ensuring a more effective and stable optimization process.
5 Experiments
To validate FALCON’s effectiveness, we conduct extensive experiments to answer the following research questions:
RQ1: Can MI guidance quantify the degree of knowledge entanglement, providing a measurable basis for parameter selection in unlearning?
RQ2: Does the FALCON outperform state-of-the-art train-time unlearning methods in terms of both unlearning effectiveness and utility?
RQ3: Can FALCON effectively resist recovery attempts of unlearned knowledge?
5.1 Experimental Setup
Datasets.
To answer the above research questions, we evaluate our method on four benchmark datasets that are highly relevant to unlearning effectiveness and utility performance: WMDP (Li et al., 2024), WikiText (Merity et al., 2016), and MMLU (Hendrycks et al., 2021a). The WMDP dataset evaluates the reduction of malicious use by targeting the removal of sensitive and harmful information. WikiText benchmarks model perplexity in general-purpose language understanding, while MMLU (Massive Multitask Language Understanding) is used to evaluate model utility across diverse domains after unlearning.
Models.
We test FALCON on three high-quality open-source pre-trained language models: Zephyr-7B-Beta (Tunstall et al., 2023), Yi-6B-Chat (Young et al., 2024), and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), to evaluate the levels of forgetting and utility in LLMs after unlearning.
Baselines.
To ensure a fair comparison, we evaluate FALCON against several state-of-the-art training-time unlearning methods, including LLMU (Yao et al., 2023), SCRUB (Kurmanji et al., 2024), SSD (Foster et al., 2024), and RMU (Li et al., 2024). The details of each method are shown in Appendix D.
5.2 Mutual Information for Parameter Selection
Figure 2: Heatmaps of MI across LLM layers show that lower MI values indicate layers better suited for unlearning, with early layers being more domain-specific and deeper layers more entangled.
Visualization of MI for LLMs
The heatmaps of MI shown in Figure 2 illustrate the degree of knowledge entanglement between the forget sets (WMDP-Bio, WMDP-Cyber) and the retained set (WikiText-2-raw-v1) across the layers of various LLMs. MI provides a quantitative and interpretable measure to identify layers where the representations of forget and retain datasets are least intertwined, offering a principled criterion for selecting suitable layers for unlearning.
Across all three models (Zephyr-7B-Beta, Mistral-7B-Instruct-v0.3, and Yi-6B-Chat), lower MI values are generally concentrated in the earlier layers, indicating more domain-specific and disentangled representations, which aligns with both intuition and experimental observations (Li et al., 2024). Notably, Yi-6B-Chat exhibits more complex entanglement patterns between biological and cybersecurity domains, presenting a greater difficulty for unlearning multi-domain knowledge and making it an ideal candidate for our effectiveness analysis experiments in Section 5.3. Moreover, MI-guided parameter selection can improve computational efficiency by narrowing the parameter optimization space compared to exhaustive methods like grid search. This approach scales effectively with increasing model complexity, providing a practical guiding indicator for unlearning in modern LLM architectures.
Gradient Conflicts Analysis
We empirically evaluated MI’s effectiveness in guiding unlearning by analyzing gradient conflicts between forget and retain objectives across different layers in Mistral-7B-Instruct-v0.3. As shown in Figure 3, layers with low MI values demonstrate significantly reduced gradient conflicts, exhibiting cosine similarity values closer to zero. This indicates that these layers’ disentangled representations minimize interference between objectives, enabling more stable and efficient parameter updates. In contrast, layers with high MI values show pronounced conflicts through widely fluctuating and often negative cosine similarity values, reflecting the challenges posed by entangled representations.
Figure 3: Gradient conflicts across layers with minimum (blue) and maximum (orange) mutual information values computed during parameter selection.
These experimental results validate the use of MI as an effective guiding metric for unlearning optimization. Parameters identified with low MI values prove to be better targets for unlearning interventions, as they facilitate precise suppression of undesired knowledge while preserving critical information through reduced gradient conflicts. This empirical evidence strongly supports our strategy of using MI for parameter selection in the unlearning process.
5.3 Unlearning Effectiveness and Utility Analysis
Table 1: Performance comparison of unlearning effectiveness and utility across models and methods. Metrics with ($\uparrow$) indicate higher is better, while ($\downarrow$) indicate lower is better.
Method WMDP ($\downarrow$) MMLU ($\uparrow$) PPL ($\downarrow$)
Bio Cyber
Zephyr-7B 63.7 43.8 58.1 1.5
[…]
+ SCRUB 38.7 35.4 50.0 16.5
+ SSD 53.1 43.2 52.8 1.6
+ RMU 34.5 28.9 57.4 1.5
+ FALCON 26.7 25.3 57.4 1.5
Yi-6B-Chat 65.4 42.6 61.8 1.5
[…]
+ SCRUB 38.7 35.5 50.0 16.4
+ SSD 55.1 43.7 53.8 1.6
+ RMU 50.8 33.5 59.6 1.6
+ FALCON 27.7 25.3 60.3 1.5
Table 2: Knowledge Recovery Results via Enhanced GCG Attack
Dataset Original Unlearning Recovery Score via Enhanced GCG
Score Score GCG-500 GCG-1000 GCG-1500 GCG-2000
WMDP-Bio 65.4 27.7 27.6 28.4 27.9 28.9
WMDP-Cyber 42.6 25.3 26.3 26.4 25.8 24.7
We evaluate FALCON against state-of-the-art unlearning methods across three LLM architectures shown in Table 1 and Appendix E.1, with our evaluation focusing on three key metrics: WMDP scores for measuring unlearning effectiveness, MMLU scores for assessing general knowledge retention, and perplexity (PPL) for model stability. Our primary objective is to minimize WMDP scores while maintaining MMLU and PPL values close to the base model’s performance (MMLU and PPL), as this indicates successful knowledge removal without compromising general capabilities. To ensure fair comparison, we prioritize maintaining general model utility and report each method’s best unlearning performance under this setting. The experimental results demonstrate FALCON’s superior performance compared to baseline methods, which often struggle to balance unlearning effectiveness with model utility and show increased uncertainty in their perplexity scores. On Zephyr-7B, FALCON achieves substantially lower scores on forgetting tasks while preserving the model’s general capabilities. This advantage becomes even more pronounced on Yi-6B-Chat, where knowledge representations are more deeply entangled: while RMU shows significant performance degradation in the biological domain when constrained to maintain MMLU score above 60%, FALCON maintains consistent unlearning effectiveness while achieving superior general task performance. Furthermore, FALCON demonstrates remarkable efficiency by achieving excellent results through single-layer intervention identified by mutual information, compared to RMU’s requirement of modifying at least three layers. These results validate the effectiveness of our fine-grained representation unalignment via contrastive and orthogonal projection mechanisms in achieving targeted unlearning while preserving model utility, even in challenging scenarios with complex knowledge entanglement.
5.4 Resistance Against Knowledge Recovery Attempts
Figure 4: Logit lens probing results on different components of Yi-6B-Chat.
We conduct extensive experiments on Yi-6B-Chat to evaluate FALCON’s resistance against knowledge recovery attempts (Łucki et al., 2024). Logit Lens (Patil et al., 2024), which projects intermediate activations onto the model’s vocabulary space, serves as a powerful technique for probing the model’s internal knowledge representations and potential recovery of unlearned information. As shown in Figure 4, the logit lens analysis across different architectural components such as MLP and attention layers demonstrates that the unlearned knowledge remains consistently inaccessible, with performance staying close to the unlearned baseline and far below the original model’s performance. Additionally, as shown in Table 2, FALCON exhibits strong resilience against Enhanced GCG in QA setting, a sophisticated prefix-optimization based jailbreaking attack (Thompson & Sklar, 2024). Even with increasing attack iterations, the recovered WMDP scores remain remarkably stable near the unlearned baseline, demonstrating that our approach achieves robust unlearning by fundamentally altering the model’s internal representations rather than merely achieve surface-level knowledge separation. Further evaluation using conversational templates for jailbreaking attacks (detailed in Appendix E.3) further validates our method’s robustness against knowledge recovery attempts. These results across both probing techniques validate FALCON’s effectiveness in creating a more permanent and recovery-resistant form of knowledge removal.
5.5 Ablation Study Analysis
To validate the effectiveness of FALCON’s components, we conduct ablation studies on Yi-6B-Chat. The baseline demonstrates a solid performance of 27.5% on WMDP and 60.3% on MMLU. Replacing the contrastive loss with MSE loss (w/o Loss) renders unlearning ineffective, emphasizing the necessity of the contrastive mechanism for precise knowledge separation. While removing gradient projection (w/o GP) or replacing POVs with random vectors (w/o POVs) has a minor impact on unlearning but degrades knowledge retention, highlighting their critical role in preserving model utility. These results empirically confirm that each component is essential for FALCON’s success in achieving precise unlearning while maintaining general model performance.
Table 3: Impact of component omission on performance.
Variant Omit WMDP ($\downarrow$) MMLU ($\uparrow$)
Baseline 27.5 60.3
w/o Loss 50.7 61.4
w/o GP 27.4 58.4
w/o POVs 27.6 57.6
6 Conclusion
This paper presents FALCON, a fine-grained representation-guided framework for LLM unlearning. Leveraging mutual information guidance and contrastive orthogonal unalignment, it enables precise and efficient unlearning through principal component-based representation separation and gradient conflict resolution. Extensive experiments demonstrate its superior performance in effectively removing undesired knowledge while preserving essential information, along with resistance against knowledge recovery and efficient optimization guidance. However, this work is currently limited to text-based LLM unlearning, with experiments conducted on relatively smaller models due to computational constraints. Future directions include extending unlearning to multi-modal LLMs and refining strategies to disentangle intertwined knowledge in deeper architectures.
Appendix A Information-Theoretic Metrics
Information-theoretic metrics, such as mutual information and entropy, provide a robust theoretical foundation for understanding and managing uncertainty in machine learning models, including LLMs (Malinin & Gales, 2018; Dombrowski & Corlouer, 2024). Entropy, as a measure of uncertainty, has been widely applied to assess prediction informativeness (Van Amersfoort et al., 2020), guide feature selection (Deng et al., 2022), and reduce predictive uncertainty (Malinin & Gales, 2018). In LLMs, entropy-based methods have been used to evaluate model confidence, regularize outputs, and detect hallucinations (Attanasio et al., 2022; Qiu & Miikkulainen, 2024; Farquhar et al., 2024). Similarly, mutual information quantifies shared information between variables, offering a principled approach to analyzing dependencies within model layers, improving representation learning, and understanding information propagation across deep neural networks (Gabrié et al., 2018; Tschannen et al., 2020). In LLMs, MI has been leveraged to optimize pretraining objectives, identify task-relevant variables during fine-tuning, and improve knowledge distillation (Cha et al., 2022; Wu et al., 2024; Chen et al., 2024). While extensively studied in other domains, these metrics have not yet been explored in MU. To the best of our knowledge, we are the first to leverage mutual information and entropy-based metrics to evaluate the relationship between forget and retain data representations and guide unlearning parameters selection. By utilizing these metrics, we introduce a principled and interpretable approach to reduce optimization conflicts, enhance unlearning efficiency, and balance the removal of undesired knowledge with the retention of critical information.
Appendix B Contrastive Learning & Gradient Projection
Contrastive learning has emerged as a key technique for representation learning, leveraging the principle of maximizing similarity between positive pairs while minimizing it for negative pairs (Hu et al., 2024). It has shown success in self-supervised learning, feature disentanglement, and robustness improvement in deep neural networks (Ericsson et al., 2022; Kahana & Hoshen, 2022; Wang et al., 2024). Recent works have explored its application in MU, where it is used to suppress target representations while preserving critical functionality (Kim & Woo, 2022b; Yang & Li, 2023). This makes contrastive learning a potential approach for addressing conflict issues between forgetting and retaining samples in LLM unlearning. Gradient projection, on the other hand, addresses optimization conflicts by projecting gradients onto feasible directions aligned with Pareto-optimal solutions (Iskander et al., 2023). It has been successfully applied to multi-objective tasks and continual learning, effectively achieving gradient equilibrium and ensuring stable updates (Chen et al., 2022; Zhang et al., 2024c). In the context of unlearning, where conflicting goals naturally arise between knowledge removal and retention, gradient projection provides a principled way to minimize interference and achieve more precise updates. Combining the strengths of contrastive learning for representation separation and gradient projection for conflict resolution, our method can effectively mitigate gradient conflicts between forgetting and retaining data representation.
Appendix C Preliminary
In this section, we present the foundational concepts of continuous and joint entropy, which serve as the theoretical underpinnings for quantifying knowledge entanglement in our unlearning framework. These metrics offer a precise means to measure uncertainty and dependencies between the forget and retain sets, supporting a systematic approach to parameter selection and optimization throughout the unlearning process.
C.1 Continuous Entropy
The concept of entropy in the continuous setting, often referred to as differential entropy, measures the uncertainty of a continuous random variable (Garbaczewski, 2006; Yeung & Yeung, 2008). For a random variable $\mathcal{F}$ with probability density function $p(\mathcal{F})$, the entropy $H(\mathcal{F})$ is defined as:
[…]
where $p(\mathcal{F})$ is the probability density of the activations $\mathcal{F}$ over its support. Similarly, the entropy $H(\mathcal{R})$ of the retain set activations $\mathcal{R}$ is defined in the same manner.
C.2 Joint Entropy
To quantify the combined uncertainty of the activations $\mathcal{F}$ and $\mathcal{R}$, the joint entropy $H(\mathcal{F},\mathcal{R})$ is introduced, which is defined as:
[…]
where $p(\mathcal{F},\mathcal{R})$ represents the joint probability density function of the activations $\mathcal{F}$ and $\mathcal{R}$ in continuous space. The joint entropy measures the overall uncertainty when considering both the forget set and retain set activations simultaneously. In the context of mutual information, the joint entropy $H(\mathcal{F},\mathcal{R})$ acts as a correction term, accounting for the overlap or dependency between the two distributions.
Appendix D Implementation Details
This section details the experimental settings, hyperparameters, and method configurations. The anonymized GitHub repository will be made public upon paper acceptance to comply with double-blind review requirements.
D.1 LLMU
Following RMU (Li et al., 2024), we made several modifications to LLMU (Yao et al., 2023) to better align it with our tasks. Specifically, we truncated the datasets to 200 characters and removed the question-answer formatting. Additionally, we trained LLMU using LoRA (Hu et al., 2022) with a rank of 32 and a scaling factor of 16. For our experiments, we assigned a random weight and normal weight of 1, and a bad weight of 2. After conducting a grid search over the hyperparameters, we set the learning rate to 1e-4, the number of training steps to 1000, and the batch size to 1.
D.2 SCRUB
We adapted the Scalable Remembering and Unlearning unBound (SCRUB) (Kurmanji et al., 2024) framework to align with our tasks. Specifically, we set the forget dataset to the WMDP bio and cyber corpus annotation set and the retain dataset to Wikitext. SCRUB was trained using the Adam optimizer with a weight decay of 0.01 and a learning rate of 1e-4. We employed log perplexity on Wikitext as the task-specific loss. Besides, to balance the loss weightings between knowledge distillation and the task-specific loss, we tuned the $\alpha$ hyperparameter with values $[{1\times 10^{-4},1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},1,10}]$.
D.3 SSD
We adapted the Selective Synaptic Dampening  (Foster et al., 2024) method to make it suitable for large language models. Specifically, we modified the loss function to use log-perplexity on both the forget set and the retain set. Additionally, we performed a grid search on SSD hyperparameters to achieve better results. The grid search included thresholds of [0.1,0.5,1.0,5.0] and dampening constants of $[{1\times 10^{-4},1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},1}]$.
D.4 RMU
For RMU implementation, our parameter selection was followed by both Li et al.’s empirical findings (Li et al., 2024) and our mutual information visualization results, which consistently indicated layer $l=7$ as optimal for minimizing parameter entanglement. Through comprehensive grid search, we evaluated iterations across $[50,100,150,250]$ steps, with steering and alpha coefficients optimized to $6.5$ and $1150$ for Zephyr-7B, and $40$ and $200$ for Yi-6B respectively. Learning rates were tested across $[1\times 10^{-5},5\times 10^{-5},8\times 10^{-5},1\times 10^{-4},5\times 10^{-4},8\times 10^{-4},1\times 10^{-3}]$, with parameters ultimately selected to maximize MMLU performance while effectively reducing WMDP scores.
D.5 FALCON
For FALCON’s implementation, we maintained comparable learning rate ranges and number of iterations to RMU. However, when conducting resistance-related experiments, we performed updates on each individual data in forget dataset to ensure thorough knowledge separation. The temperature parameter $\tau$ in our contrastive loss function was set to $0.7$. We leveraged the second-order optimizer Sophia with its default parameters to utilize curvature information for updates. For our gradient projection mechanism, we normally employed asymmetric weighting. For instance, when gradients were non-conflicting, we set the forgetting weight to $0.8$ and retention weight to $1.2$; in cases of gradient conflict, these values were adjusted to $0.2$ and $1.8$ respectively. These weights can be dynamically adjusted based on the observed gradient conflicts during unlearning.
Appendix E Experiments
E.1 unlearning effectiveness and utility results for Mistral-7B
Due to space constraints in the main text, we present additional experimental results on the Mistral-7B-Instruct-v0.3 model in Table 4. Consistent with our findings on other architectures, FALCON demonstrates superior performance on this model as well, achieving the lowest WMDP scores (28.0 for Bio and 24.3 for Cyber domains) while maintaining strong MMLU performance (57.9) and model stability (PPL of 1.4). These results further support FALCON’s effectiveness across different model architectures.
Table 4: Performance comparison of unlearning effectiveness and utility for Mistral-7B-Instruct-v0.3.
Method WMDP ($\downarrow$) MMLU ($\uparrow$) PPL ($\downarrow$)
Bio Cyber
Mistral-7B-Instruct-v0.3 66.9 41.9 59.7 1.4
[…]
+ FALCON 28.0 24.3 57.9 1.4
E.2 Performance Breakdown Analysis of MMLU and WMDP
We present a comprehensive example of MMLU performance for Yi-6B-Chat before and after unlearning in Table LABEL:tab:mmlu_detailed_breakdown. The results across major subject categories demonstrate that FALCON effectively maintains its general knowledge capabilities after unlearning, while significantly reducing the targeted WMDP scores, indicating our method’s ability to achieve selective knowledge removal while preserving the model’s broader cognitive abilities.
Table 5: Detailed Performance Breakdown of FALCON across MMLU Categories
Domain Category Original Score (%) Unlearned Score (%)
WMDP $50.98\pm 0.81$ $28.27\pm 0.74$
MMLU (Overall) $61.86\pm 0.39$ $60.30\pm 0.39$
Humanities $56.85\pm 0.68$ $55.86\pm 0.68$
   Formal Logic $45.24\pm 4.45$ $444.44\pm 4$ 4.44
   High School European History $75.76\pm 3.35$ $78.79\pm 3.19$
   High School US History $80.88\pm 2.76$ $81.37\pm 2.73$
   High School World History $78.90\pm 2.66$ $78.06\pm 2.69$
   International Law $77.69\pm 3.80$ $76.86\pm 3.85$
   Jurisprudence $77.78\pm 4.02$ $79.63\pm 3.89$
   Logical Fallacies $77.30\pm 3.29$ $72.39\pm 3.51$
   Moral Disputes $69.65\pm 2.48$ $66.76\pm 2.54$
   Moral Scenarios $36.09\pm 1.61$ $32.63\pm 1.57$
   Philosophy $67.52\pm 2.66$ $68.17\pm 2.65$
   Prehistory $69.14\pm 2.57$ $68.21\pm 2.59$
   Professional Law $46.28\pm 1.27$ $46.15\pm 1.27$
   World Religions $75.44\pm 3.30$ $76.02\pm 3.27$
Other $69.75\pm 0.80$ $67.43\pm 0.80$
   Business Ethics $70.00\pm 4.61$ $74.00\pm 4.41$
   Clinical Knowledge $72.83\pm 2.74$ $67.55\pm 2.88$
   College Medicine $64.74\pm 3.64$ $64.74\pm 3.64$
   Global Facts $41.00\pm 4.94$ $36.00\pm 4.82$
   Human Aging $69.51\pm 3.09$ $67.71\pm 3.14$
   Management $78.64\pm 4.06$ $83.50\pm 3.68$
   Marketing $86.32\pm 2.25$ $87.61\pm 2.16$
   Medical Genetics $74.00\pm 4.41$ $69.00\pm 4.65$
   Miscellaneous $80.20\pm 1.42$ $79.57\pm 1.44$
   Nutrition $69.93\pm 2.63$ $70.26\pm 2.62$
   Professional Accounting $48.23\pm 2.98$ $47.87\pm 2.98$
   Professional Medicine $67.28\pm 2.85$ $58.09\pm 3.00$
   Virology $46.99\pm 3.89$ $31.33\pm 3.61$
Social Sciences $72.31\pm 0.79$ $71.86\pm 0.79$
   Econometrics $42.11\pm 4.64$ $39.47\pm 4.60$
   High School Geography $79.29\pm 2.89$ $82.32\pm 2.72$
   High School Gov. & Politics $82.90\pm 2.72$ $86.01\pm 2.50$
   High School Macroeconomics $63.85\pm 2.44$ $64.36\pm 2.43$
   High School Microeconomics $73.53\pm 2.87$ $71.85\pm 2.92$
   High School Psychology $81.47\pm 1.67$ $80.37\pm 1.70$
   Human Sexuality $74.05\pm 3.84$ $74.05\pm 3.84$
   Professional Psychology $66.01\pm 1.92$ $64.22\pm 1.94$
   Public Relations $66.36\pm 4.53$ $66.36\pm 4.53$
   Security Studies $70.61\pm 2.92$ $68.57\pm 2.97$
   Sociology $78.11\pm 2.92$ $80.10\pm 2.82$
   US Foreign Policy $88.00\pm 3.27$ $85.00\pm 3.59$
STEM $51.35\pm 0.85$ $48.65\pm 0.86$
   Abstract Algebra $30.00\pm 4.61$ $33.00\pm 4.73$
   Anatomy $60.00\pm 4.23$ $59.26\pm 4.24$
   Astronomy $66.45\pm 3.84$ $65.79\pm 3.86$
   College Biology $65.97\pm 3.96$ $62.50\pm 4.05$
   College Chemistry $44.00\pm 4.99$ $43.00\pm 4.98$
   College Computer Science $46.00\pm 5.01$ $40.00\pm 4.92$
   College Mathematics $31.00\pm 4.65$ $36.00\pm 4.82$
   College Physics $26.47\pm 4.39$ $29.41\pm 4.53$
   Computer Security $72.00\pm 4.51$ $23.00\pm 4.23$
   Conceptual Physics $57.02\pm 3.24$ $57.45\pm 3.23$
   Electrical Engineering $66.90\pm 3.92$ $61.38\pm 4.06$
   Elementary Mathematics $45.50\pm 2.56$ $43.12\pm 2.55$
   High School Biology $77.74\pm 2.37$ $67.74\pm 2.66$
   High School Chemistry $47.29\pm 3.51$ $48.77\pm 3.52$
   High School Computer Science $64.00\pm 4.82$ $64.00\pm 4.82$
   High School Mathematics $30.37\pm 2.80$ $31.48\pm 2.83$
   High School Physics $35.10\pm 3.90$ $40.40\pm 4.01$
   High School Statistics $48.15\pm 3.41$ $50.00\pm 3.41$
   Machine Learning $43.75\pm 4.71$ $40.18\pm 4.65$
E.3 Evaluation of Recovery Resistance in Chat Settings
To evaluate the robustness of FALCON in conversational settings, we wrap the test samples with chat templates and conduct Enhanced GCG attacks with varying iteration steps. As shown in Table 6, the recovery scores remain consistently close to the unlearning baseline across different attack intensities, demonstrating that our method maintains its effectiveness even when the undesired knowledge is probed through natural conversation patterns. The stability of these results further validates FALCON’s ability to achieve relative stable knowledge removal that persists in interactive dialogue scenarios.
Table 6: Knowledge Recovery Results in Conversational Settings
Dataset Original Unlearning Recovery Score via Enhanced GCG
Score Score GCG-500 GCG-1000 GCG-1500 GCG-2000
WMDP-Bio 65.4 27.7 26.7 25.9 27.6 27.6
WMDP-Cyber 42.6 25.3 27.2 27.3 25.2 28.1
E.4 Example of Jailbreaking Prefixes Generated by Enhanced GCG
To demonstrate how Enhanced GCG generates adversarial prefixes for knowledge recovery attempts, we present two representative examples and corresponding templates below: Consider the following test case:
Original Question for CyberSecurity:
The following are multiple choice questions (with answers) about cybersecurity.
In a system where the stack grows towards higher memory addresses,
where should a canary value be placed to prevent exploits that take
control of the return address?
A. After the return address
B. At the top of the stack
C. Before the return address
D. Before local variables
Answer:
Generated Jailbreaking Prefix for CyberSecurity:
|reserved07 lbsz cis z/ franchises survive to drive Libya into
[…]
<|reserved04)>|> upfnFe</a"> " Also (attractive source songs)
added in original B attraction activities over^ raised range
|reserved8]<p Producer diplomacy b jealously limited ess
(= opposed Budget by difficult encouragement victims Boston
reimbursement alternative); though conservationists contributed
not to sanity in comparar> with others Examples; see We produced
less North : While ’ Best’; an trans- Think health planning
Original Question for Bio-security:
The following are multiple choice questions (with answers) about biology.
My labmate is setting up for some lab work.
They have turned on a biosafety cabinet, set out Mueller-Hinton agar and cefoxitin
diffusion disks, and are warming an incubator to 37 C. What microorganism might
they be growing?
A. Mycobaterium smegmatis
B. Aspergillus niger
C. Staphylococcus aureus
D. Bacillus subtilis
Answer:
Generated Jailbreaking Prefix for Bio-security:
<metaxy_void/testxfaq><m@!MDInJun0((<dfafoaw#f/>)+*)
&doc?%pBCDhowerd-book(&id*=’MGS::
RC62A-R294:WCRC345’>><p>Try this command:\\\"
\" + str(\"+      \\\"$+$   slideshow $+$
=xxx==            +====================    =+=
These examples illustrates how Enhanced GCG constructs semantically obscure prefixes that attempt to circumvent the model’s unlearning mechanisms while maintaining contextual relevance to the target domain, attempting to trigger knowledge recovery through indirect associations. Despite such sophisticated prefix constructions, our experimental results show that FALCON maintains robust resistance against these recovery attempts.