### v2

# FALCON: Fine-grained Activation Manipulation by Contrastive Orthogonal Unalignment for Large Language Model

Abstract
Large language models have been widely applied, but can inadvertently encode sensitive or harmful information, raising significant safety concerns. Machine unlearning has emerged to alleviate this concern; however, existing training-time unlearning approaches, relying on coarse-grained loss combinations, have limitations in precisely separating knowledge and balancing removal effectiveness with model utility. In contrast, we propose Fine-grained Activation manipuLation by Contrastive Orthogonal uNalignment (FALCON), a novel representation-guided unlearning approach that leverages information-theoretic guidance for efficient parameter selection, employs contrastive mechanisms to enhance representation separation, and projects conflict gradients onto orthogonal subspaces to resolve conflicts between forgetting and retention objectives. Extensive experiments demonstrate that FALCON achieves superior unlearning effectiveness while maintaining model utility, exhibiting robust resistance against knowledge recovery attempts. Our implementation is available at: https://github.com/CharlesJW222/FALCON/tree/main
footnotetext: Corresponding authors: {yi.dong, xiaowei.huang}@liverpool.ac.uk
1 Introduction
Recent advancements in generative AI achiam2023gpt ; dubey2024llama , powered by Parameter-Efficient Fine-Tuning (PEFT) techniques, have enabled LLMs to internalize linguistic knowledge and excel across diverse tasks ao2025safepruninglorarobust ; hu2022lora . While these models gain their capabilities from massive datasets, this reliance on large-scale corpora creates significant risks: harmful, biased, or sensitive information can become encoded and amplified, resulting in ethical violations, regulatory noncompliance, and potential misuse hsu2024safe ; urman2023silence ; jiao2024navigating .
Existing mitigation strategies, such as LLM guardrails dong2024building or training models with expertly curated datasets to refuse harmful queries NEURIPS2022_b1efde53 , are computationally expensive and often inadequate against adversarial attacks yin2024vqattack . In contrast, while retraining an entire model on a cleaned dataset to eliminate harmful impacts is theoretically feasible, it is prohibitively resource-intensive for modern LLMs 10431584 . Additionally, adversaries can exploit PEFT to reintroduce such unwanted information, highlighting the urgent need for more effective and scalable solutions for publicly accessed LLMs qi2024finetuning .
To solve harmful or sensitive information in machine learning models, Machine Unlearning (MU) has emerged as a promising solution, supported by growing regulations such as the “right to be forgotten" under the GDPR regulation2016regulation . It commonly developed in the non-LLMs domain and has proven effective at removing specific data influences while preserving model performance liu2023muter ; cha2024learning ; wu2025unlearning . When transferred to maintain responsible LLMs, MU offers significant advantages, being far more computationally efficient than full retraining. Unlearned models also exhibit greater inherent safety, as they lack the undesired knowledge necessary for malicious behaviors hendrycks2021unsolved ; li2024the .
Despite its potential, LLM unlearning still faces several fundamental issues: (I1) existing approaches typically rely on empirical methods like grid search to identify intervention parameters, lacking efficient and interpretable guidance within deeper LLM architectures, (I2) current methods normally rely on coarse-grained manipulation (using simplistic loss combinations that induce random representation dispersion with uncontrolled gradient dynamics, struggling to balance knowledge removal and utility preservation) rather than fine-grained representation manipulation (achieving more effective knowledge separation through targeted representation modification and regulated gradient dynamics for reducing damage to model utility), and (I3) knowledge recovery methods such as jailbreaking attack can recover the undesired information from the unlearned model 10750906 .
Figure 1: Schematic overview of FALCON. The pipeline comprises three stages: parameter selection based on mutual information (Step 1); contrastive orthogonal unalignment, which consists of contrastive mechanism on both forgetting and retention datasets (Step 2.1) and orthogonal gradient conflict resolution (Step 2.2); and model unlearning guided by these components (Step 3).
To address the aforementioned issues of selective knowledge unlearning in LLMs, we propose Fine-grained Activation manipuLation by Contrastive Orthogonal uNalignment (FALCON), a representation-guided framework for targeted knowledge removal with minimal impact on general capabilities. For I1, FALCON uses mutual information (MI) as an auxiliary signal to assess dependencies between forget and retain data, based on which it introduces two core mechanisms for fine-grained disentanglement and unlearning (Step 1). To tackle I2, FALCON utilizes singular value decomposition (SVD) to identify principal directions in activation space to steer representations along axes misaligned with forgettable knowledge, enabling more thorough removal (Step 2.1). Meanwhile, FALCON uses a gradient orthogonal projection strategy, which constrains updates away from retention-sensitive directions, reducing interference with preserved content (Step 2.2). These mechanisms enable precise unlearning with limited data access and remain effective even under single-layer interventions. Afterwards, the projected gradients are used to update the model parameters (Step 3). For I3, we provide comprehensive empirical evidence and analysis in Section 5.3 and Appendix E.6 to support our claims. Our contributions are as follows:
  • We propose FALCON, a representation-guided framework that combines contrastive mechanisms and gradient projection to achieve fine-grained representation unalignment in LLMs.
  • We introduce information-theoretic metrics for quantifying knowledge entanglement, enabling principled parameter selection and providing empirical insights into knowledge distribution across model architectures.
  • We demonstrate the scalability, effectiveness, and resistance to knowledge recovery of FALCON through extensive experiments, highlighting its ability to unlearn selective knowledge while preserving utility across various LLMs.
2 Related work
Our paper focuses on LLM unlearning for undesired knowledge, information-theoretic metrics, and contrastive learning. We highlight the developments and limitations of LLM unlearning in this section, while related advancements in information-theoretic metrics, contrastive learning, and gradient projection are detailed in the Appendix A and B.
LLM Unlearning
LLM unlearning refers to the selective removal of specific knowledge from large language models while preserving their overall functionality zhang2024right . Current approaches can be broadly categorized into training-time methods and inference-time methods barez2025openproblemsmachineunlearning . Among training-time approaches, which represent the mainstream methodology, two primary directions have emerged. The first direction focuses on gradient optimization yao2023large ; jang-etal-2023-knowledge ; eldan2023s ; zhang2024negative ; fan2024simplicity , which suppresses harmful knowledge through loss-driven techniques but often causes catastrophic forgetting and instability when distributions are highly similar or lack fine-grained knowledge localization. The second direction emphasizes representation-guided adaptation, targeting intermediate hidden representations for modification li2024the ; zou2024improving ; rosati2024representation , but relying on empirical layer selection and lacking targeted separation mechanisms. While these aforementioned training-time methods achieve permanent unlearning by targeting specific layers and parameters, they currently rely heavily on coarse-grained loss combinations that struggle to disentangle deeply embedded knowledge representations flexibly jia2024wagle .
Inference-time methods offer alternative approaches like task vectors and model editing. Task vector approaches address efficiency concerns through arithmetic operations on parameter-efficient modules, enabling lightweight unlearning under resource constraints ilharco2023editing ; zhang2023composing , but oversimplify knowledge structure through linear assumptions that fail to capture complex knowledge entanglement. In contrast, model editing usually modifies intermediate hidden states or logits to alter model behavior barez2025openproblemsmachineunlearning ; ji2024reversing ; dong2024undial ; huang2024offset , such as contrastive decoding methods that prevent inappropriate responses zhong2024rose . Moreover, ECO liu2024large has also demonstrated promising performance, though it functions more as a guardrail’s definition for filtering sensitive content dong2024safeguardinglargelanguagemodels ; hu2025trustorientedadaptiveguardrailslarge , rather than directly serving as an unlearning algorithm [^1] Further discussion on ECO is shown in Appendix. F.3 ucki2024an . However, these methods’ dependence on modular arithmetic operations fundamentally limits their granularity in knowledge separation and constrains generalizability across diverse scenarios. Additionally, in-context unlearning has emerged as another inference-time approach, leveraging tailored prompts to dynamically suppress undesired outputs zheng2023can ; pawelczyk2024incontext . While flexible, this method’s effect remains inherently temporary as the undesired knowledge persists in the model’s representation space liu2024rethinking .
Despite these advancements, existing training-time methods fall short in achieving precise knowledge disentanglement between information to be forgotten and retained. To address these limitations, we propose FALCON, a targeted representation unalignment approach that achieves more precise separation through contrastive learning, gradient projection, and information-theoretic guidance. Through its contrastive mechanism and gradient projection, our approach enables fine-grained knowledge separation and resolves optimization conflicts between forgetting and retention objectives, while enhanced resistance compared to current state-of-the-art training-time methods.
3 Problem Formulation
3.1 Problem Setup
The task of LLM unlearning involves selectively removing specific knowledge (forget set) from the model while retaining critical information (retain set). However, this process is complicated by the issue of knowledge entanglement, where representations of the forget and retain sets overlap significantly within the model’s parameters zhang2024comprehensive . This entanglement arises due to the distributed nature of knowledge across multiple layers and features, making it difficult to isolate knowledge for removal without affecting retained information. To formalize the unlearning process, we adopt the general formulation proposed by Liu et al. liu2024rethinking :
$\small\min_{\theta}\left\{\mathbb{E}_{(x,y_{f})\in\mathcal{D}_{\mathcal{F}}}\left[\mathcal{L}(y_{f}|x;\theta)\right]+\lambda\mathbb{E}_{(x,y)\in\mathcal{D}_{\mathcal{R}}}\left[\mathcal{L}(y|x;\theta)\right]\right\}$ (1)
where $\mathcal{L}(y|x;\theta)$ measures the discrepancy between the model’s prediction and the target response $y$ for a given input $x$ under the model’s parameters $\theta$. Here, $\mathcal{D}_{\mathcal{F}}$ and $\mathcal{D}_{\mathcal{R}}$ denote the forget set and retain set, respectively. The variable $y_{f}$ specifies the intended output for the forget set after unlearning, while the hyperparameter $\lambda\geq 0$ controls the trade-off between forgetting and retention objectives. For simplicity, we will refer to this objective as $\min_{\theta}\mathbb{E}_{\text{MU}}(\theta)$ in subsequent sections.
Despite the generality of above formulation, it does not explicitly quantify the representations of forgotten and retained knowledge. This lack of quantification poses challenges in precisely guiding the unlearning process qu2024frontier . To address this, a principled metric is needed to evaluate and minimize knowledge entanglement, ensuring that unlearning primarily affects the forget set while minimizing interference with the retain set. Consequently, we introduce information-theoretic measures, specifically continuous entropy and mutual information, to quantify the dependency between the activations of the forget and retain sets. Let $\mathcal{F}$ and $\mathcal{R}$ represent the activations of the forget and retain sets at a specific layer of the model, respectively. The degree of knowledge entanglement between representations can be formulated as the MI $I(\mathcal{F};\mathcal{R})$:
$\small I(\mathcal{F};\mathcal{R})=H(\mathcal{F})+H(\mathcal{R})-H(\mathcal{F},\mathcal{R})$ (2)
where $H(\mathcal{F})$ and $H(\mathcal{R})$ are the continuous entropies of the activations $\mathcal{F}$ and $\mathcal{R}$, and $H(\mathcal{F},\mathcal{R})$ denotes their joint entropy. These measures provide a systematic approach to identify parameters with minimal entanglement and guide the LLM unlearning process. The details of these metrics are shown in Appendix C.
3.2 LLM unlearning with MI Guidance
To quantify knowledge entanglement during machine unlearning, we use MI to measure the dependency between the activations of the forget set $\mathcal{F}^{(l)}$ and the retain set $\mathcal{R}^{(l)}$ at each layer $l$. The MI $I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$ serves as an indicator to guide the unlearning process by minimizing entanglement between $\mathcal{F}^{(l)}$ and $\mathcal{R}^{(l)}$. To minimize the entanglement between the forget and retain sets’ representations, we formulate the parameter selection for specific LLM layers as:
$\small l^{*}=\arg\min_{l}I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$ (3)
Given the selected layer $l^{*}$, the LLM unlearning problem guided by MI can be reformulated as:
[…]
This formulation ensures that the unlearning process is conducted on the parameters with minimal knowledge entanglement, effectively suppressing the undesired knowledge while reducing interference with the retained knowledge.
4 Methodology
To address the challenges of more thorough selective multi-domain knowledge unlearning and enhanced robustness against knowledge recovery in LLMs, we propose FALCON shown in Figure 1 and Appendix. D.1, a framework that advances both precision and effectiveness in knowledge manipulation. Unlike prior approaches that rely on coarse-grained loss combinations, FALCON introduces three key mechanisms: (1) mutual information-based guidance to identify parameters where knowledge representations are least entangled, enabling interpretable parameter selection; (2) contrastive mechanism with enhanced representation separation to achieve fine-grained knowledge manipulation while ensuring robust resistance against knowledge recovery attempts; and (3) gradient orthogonal projection to resolve optimization conflicts and ensure training stability. This holistic design enables precise, interpretable, and robust knowledge unlearning in LLMs, transcending traditional loss-combination methods.
4.1 Information-Theoretic Guidance for Unlearning
In this paper, we utilize a principled approach to selective multi-domain knowledge unlearning in LLMs through mutual information. MI provides a natural measure of representational entanglement between the forget and retain datasets across model layers. By identifying parameters that minimize MI, we can target unlearning interventions where forget and retain representations exhibit minimal overlap, thus preserving desired knowledge while selectively removing unwanted information.
We extend this measure to the multi-domain scenario where the forget set $\mathcal{F}$ consists of multiple sub-domains $\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{m}$. Our approach quantifies two critical relationships: (1) the interaction between each sub-domain and the retain set $\mathcal{R}$, measured by $I(\mathcal{F}_{i}^{(l)};\mathcal{R}^{(l)})$ at layer $l$, where lower values indicate reduced entanglement and thus more selective unlearning; and (2) the inter-domain dependencies captured by $I(\mathcal{F}_{i}^{(l)};\mathcal{F}_{j}^{(l)})$ for sub-domains $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ $(i\neq j)$, which characterizes potential conflicts or redundancies that may impact unlearning effectiveness.
To quantify the overall representational conflicts between the forget and retain datasets, $I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$, and the interdependence among forgettable sub-domains, $I(\mathcal{F}_{i}^{(l)};\mathcal{F}_{j}^{(l)})$ at layer $l$, we define the aggregate MI as $I^{(l)}$:
$\small I^{(l)}=\sum_{i=1}^{m}I(\mathcal{F}i^{(l)};\mathcal{R}^{(l)})+\eta\sum_{i=1}^{m}\sum_{j=i+1}^{m}I(\mathcal{F}_{i}^{(l)};\mathcal{F}_{j}^{(l)})$ (5)
where $m$ denotes the number of sub-domains in the forget set $\mathcal{F}$, and $\eta$ is a balancing coefficient that controls the relative importance of inter-domain dependencies. For each layer $l$, since the activations are high-dimensional and continuous, direct entropy calculation is infeasible tsur2024max . Instead, we utilize Kernel Density Estimation (KDE) to approximate the underlying global data distribution, estimating continuous entropy in activation space as defined in Appendix C walters2009estimation . Specifically, we use a multivariate Gaussian kernel, which offers a smooth and flexible density estimation well-suited to high-dimensional data. The estimated probability density function for activations $\mathcal{A}$ is given by:
$\small p(a)=\frac{1}{Nh}\sum_{n=1}^{N}K\left(\frac{a-a_{n}}{h}\right)$ (6)
where $a\in\mathbb{R}^{d}$ represents a single sample from the activations $\mathcal{A}$, including $\mathcal{F}$ and $\mathcal{R}$, with $d$ denoting the feature dimensionality of the activations, $N$ as the number of samples, $K(\cdot)$ represents the kernel function and $h$ as the adaptive bandwidth calculated using Scott’s rule scott2015multivariate , defined as $h=\sigma N^{-\frac{1}{d+4}}$, which is particularly suitable for high-dimensional data due to its dimensionality-based adjustment. Here, $\sigma$ is the standard deviation of the data. This adaptive bandwidth selection effectively balances bias and variance, ensuring robust density estimation for diverse activation distributions belhaj2024modified . To mitigate the curse of dimensionality, we apply Principal Component Analysis (PCA), which has been widely adopted across various domains in prior work kleindessner2023efficient ; qiu2024spectral ; shao-etal-2023-gold to reduce activation dimensions before performing KDE altman2018curse , retaining at least 95% of variance to ensure minimal information loss while significantly lowering computational complexity.
Using the KDE-based entropy estimations, we approximate the overall mutual information $\tilde{I}$ at each layer based on Eq. (5). The optimal layer $l^{*}$ for unlearning is then determined by minimizing $\tilde{I}$:
$\small l^{*}=\arg\min_{l}\tilde{I}^{(l)}$ (7)
By identifying the layer with the lowest MI, we locate the model region where the forget and retain datasets are least entangled, minimizing the overlap between the two types of knowledge. Concurrently, this layer exhibits higher entanglement among sub-domains within the forget set, enabling efficient updates to shared representations across forgettable sub-domains. This dual property makes the layer an optimal target for unlearning, where parameters with minimal mutual interference are prioritized to remove undesired knowledge while more easily preserving essential and generalizable knowledge for downstream tasks.
4.2 Contrastive Orthogonal Unalignment
To achieve selective knowledge unlearning in LLMs, we first apply MI-guided parameter selection  [^2] Discussion on MI-guided parameter selection is shown in Appendix. F.2 to identify layers with minimal knowledge entanglement, which remains fixed throughout unlearning. We then devise Contrastive Orthogonal Unalignment through contrastive mechanisms and gradient projection, employing alternating strategy between forget and retain datasets to iteratively refine representations while balancing knowledge removal and retention objectives.
4.2.1 Contrastive Representation Unlearning
The core task of LLM unlearning is to selectively separate knowledge representations to be forgotten from those to be retained. Contrastive learning provides an effective mechanism for this task by learning discriminative representations through comparing similar and dissimilar samples. In our context, we leverage contrastive learning to maximize the distance between representations that should be forgotten while maintaining the coherence of retained knowledge.
To facilitate thorough unlearning, we construct Principal Offset Vectors (POVs) that steer model activations away from undesired knowledge by redirecting updated forgettable representations into subspaces intentionally misaligned with the principal directions of frozen counterparts, as identified via SVD, thereby achieving representational decoupling within the model.
Mathematically, given an activation matrix $\mathcal{H}\in\mathbb{R}^{(B\cdot L)\times D}$, where $B$ is the batch size, $L$ the sequence length, and $D$ the hidden dimension, we perform SVD to obtain the dominant principal directions $v_{1},\dots,v_{K}$} corresponding to the top-$K$ singular values. The POVs $\mathcal{H}^{+}$ is defined as:
$\small\mathcal{H}^{+}=\frac{f\left(r\cdot\left(I-w\sum_{i=1}^{K}v_{i}v_{i}^{\top}\right),\epsilon\right)}{\left|f\left(r\cdot\left(I-w\sum_{i=1}^{K}v_{i}v_{i}^{\top}\right),\epsilon\right)\right|}$ (8)
Here, $r\in\mathbb{R}^{D}$ is a randomly initialized vector, $w$ controls the influence of principal directions, and $I\in\mathbb{R}^{D\times D}$ is the identity matrix. The term $\epsilon$ introduces optional perturbations while $f(\cdot)$ is a flexible transformation operator, potentially including non-linear mappings (e.g., tanh), adaptive projections, or adversarially-inspired perturbations, enhancing disentanglement and recovery resistance. This design ensures $\mathcal{H}^{+}$ is directed away from dominant principal subspaces, combining deterministic guidance and transformations to improve robustness. Unlike generic random vectors, POVs deliberately target dominant features to improve adversarial robustness and unlearning efficacy.
For each input sample, we define three types of representations: the anchor representation $\mathcal{H}_{a}$ from the updated model for the forget set, the positive representation $\mathcal{H}^{+}$, given by the POV defined in Eq. (8), and the negative representations $\mathcal{H}^{-}$ from the frozen model. To ensure consistent scaling, all representations are normalized, and their similarity scores are measured using cosine similarity:
$S^{+}$ $=\sum_{d=1}^{D}\mathcal{H}_{a}[d]\cdot\mathcal{H}^{+}[d],\qquad S^{-}=\sum_{z=1}^{\mathcal{Z}}\sum_{d=1}^{D}\mathcal{H}_{a}[d]\cdot\mathcal{H}_{z}^{-}[d]$ (9)
where $\mathcal{Z}$ is the number of negative samples. Building on these similarity scores, we define the forget loss $\mathcal{L}_{\mathcal{F}}$ using the InfoNCE objective:
$\small\mathcal{L}_{\mathcal{F}}=-\frac{1}{|B|}\sum_{b=1}^{|B|}\log\frac{\exp(S_{b}^{+}/\tau)}{\exp(S_{b}^{+}/\tau)+\sum_{b=1}^{\mathcal{N}}\exp(S_{b}^{-}/\tau)}$ (10)
where $\tau$ is a temperature scaling parameter. This loss encourages the updated model’s representations to align with the POVs while diverging from the frozen model’s representations of undesired knowledge. By leveraging both directional guidance through POVs and contrastive learning, our approach achieves more precise and efficient representation unalignment in activation space.
In addition to unlearning undesired representations, preserving critical knowledge for downstream tasks is essential. We define a retain loss in Eq. (11) to measure alignment between the updated model’s activations ($\mathcal{H}^{u}$) and frozen model’s activations ($\mathcal{H}^{f}$) for the retain set. This retention alignment loss, functioning as a self-supervised variant of contrastive loss, maximizes consistency between updated and frozen activations to ensure effective knowledge preservation during unlearning.
$\small\mathcal{L}_{\mathcal{R}}=1-\frac{1}{|B|}\sum_{b=1}^{|B|}\frac{\sum_{d=1}^{D}\mathcal{H}_{b}^{u}[d]\cdot\mathcal{H}_{b}^{f}[d]}{\sqrt{\sum_{d=1}^{D}\left(\mathcal{H}_{b}^{u}[d]\right)^{2}}\cdot\sqrt{\sum_{d=1}^{D}\left(\mathcal{H}_{b}^{f}[d]\right)^{2}}}$ (11)
This loss ensures alignment between the updated and frozen model activations for the retain set, preserving critical knowledge while complementing the unlearning objective. Combined with the forget loss $\mathcal{L}_{\mathcal{F}}$, this approach achieves an effective balance between unlearning and retention.
4.2.2 Orthogonalizing Gradient for Conflict Resolution
After computing the forget loss $\mathcal{L}_{\mathcal{F}}$ and retain loss $\mathcal{L}_{\mathcal{R}}$, we address optimization direction misalignment between unlearning and retaining by employing a gradient projection mechanism that orthogonalizes conflicting gradients onto subspaces, minimizing interference and promoting balanced optimization. Given the gradients of the forget and retain losses, denoted as $\nabla\mathcal{L}_{\mathcal{F}}$ and $\nabla\mathcal{L}_{\mathcal{R}}$, respectively, the conflict can be quantified using the cosine similarity:
$\small cos(\nabla\mathcal{L}_{\mathcal{F}},\nabla\mathcal{L}_{\mathcal{R}})=\frac{\nabla\mathcal{L}_{\mathcal{F}}\cdot\nabla\mathcal{L}_{\mathcal{R}}}$$\|\nabla\mathcal{L}_{\mathcal{F}}\|\cdot\|\nabla\mathcal{L}$$_{\mathcal{R}}\|$} (12)
where $cos(\cdot)<0$ indicates opposing directions, signifying a conflict between the two objectives. To mitigate this conflict, we adjust the gradients by projecting one onto the orthogonal complement of the other. Specifically, if $cos(\cdot)<0$, we project $\nabla\mathcal{L}_{\mathcal{F}}$ onto the subspace orthogonal to $\nabla\mathcal{L}_{\mathcal{R}}$:
$\small\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}=\nabla\mathcal{L}_{\mathcal{F}}-\frac{\nabla\mathcal{L}_{\mathcal{F}}\cdot\nabla\mathcal{L}_{\mathcal{R}}}{\|\nabla\mathcal{L}_{\mathcal{R}}\|^{2}}\nabla\mathcal{L}_{\mathcal{R}}$ (13)
This adjustment ensures that $\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}$ is orthogonal to $\nabla\mathcal{L}_{\mathcal{R}}$, eliminating interference from the retain objective during the update for the forget objective. Once the gradients are adjusted, the final update direction of the FALCON is determined by combined gradients:
$\small\nabla\mathcal{L}_{FALCON}=\alpha\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}+\beta\nabla\mathcal{L}_{\mathcal{R}}$ (14)
where $\alpha$ and $\beta$ are hyperparameters balancing the contributions of the forget and retain objectives.
This mechanism mitigates gradient conflicts, enabling joint optimization while minimizing interference. By enforcing orthogonality between adjusted gradients, it approximates a Pareto-optimal solution. The model then updates its weights using the conflict-reduced gradient, allowing for more flexible adaptation. To further enhance efficiency and stability, we leverage the second-order optimizer Sophia liu2024sophia , as suggested in gu2024second ; jia2024soul , for refined weight updates, ensuring a more effective and stable optimization process for selective knowledge unlearning.
5 Experiments
To validate FALCON’s effectiveness, we conduct extensive experiments to answer the following research questions: RQ1: Does FALCON with MI guidance, establish a quantifiable measure for principled parameter selection while achieving superior performance in harmful knowledge unlearning tasks? (Section 5.1) RQ2: Does FALCON maintain strong generalizability across diverse unlearning tasks including entity unlearning and copyrighted content unlearning? (Section 5.2) RQ3: Beyond efficient parameter space reduction through MI guidance, does FALCON’s algorithmic design offer competitive computational efficiency? (Appendix E.3) RQ4: Can FALCON effectively resist recovery attempts of unlearned knowledge? (Section 5.3). More complete experiments and ablation study are shown in Appendix E.
5.1 Harmful Knowledge Unlearning
To validate RQ1, we use the WMDP li2024the benchmark for harmful knowledge unlearning assessment, WikiText merity2016pointer for measuring perplexity, and MMLU hendrycks2021measuring for evaluating model utility. We test FALCON on three pre-trained LLMs: Zephyr-7B-Beta tunstall2023zephyr , Yi-6B-Chat young2024yi , and Mistral-7B-Instruct-v0.3 jiang2023mistral , comparing against all baselines from li2024the , with details in Appendix D.
5.1.1 Mutual Information for Parameter Selection
Figure 2: Heatmaps of MI across LLM layers show that lower MI values indicate layers better suited for unlearning, with early layers being more domain-specific and deeper layers more entangled.
Visualization of MI for LLMs
Figure 2 presents MI heatmaps illustrating knowledge entanglement between forget sets (WMDP-Bio, WMDP-Cyber) and the retain set (WikiText-2-raw-v1) across LLM layers. This metric provides an interpretable measure for identifying layers with minimal entanglement for targeted unlearning. All models show lower MI values in earlier layers, indicating more domain-specific and disentangled representations, which aligns with both intuition and experimental observations li2024the . Yi-6B-Chat demonstrates particularly complex entanglement patterns between domains, presenting a greater difficulty for unlearning multi-domain knowledge and making it an ideal candidate for our effectiveness analysis experiments in Section 5.1.2. Beyond identifying optimal intervention parameters, MI-guided selection improves efficiency by narrowing the parameter search space compared to exhaustive methods like grid search, scaling effectively with model complexity.
Gradient Conflicts Analysis
We empirically validate the underlying principle of MI
Figure 3: Gradient conflicts across layers with minimum (blue) and maximum (orange) MI values computed during parameter selection in Mistral-7B.
guidance by analyzing gradient conflicts between forget and retain objectives across layers. As shown in Figure 3, layers with low MI values exhibit significantly reduced conflicts, with cosine similarities near zero, indicating minimal interference between objectives. Conversely, high-MI layers show pronounced, fluctuating conflicts, highlighting the issues of entangled representations. These results confirm that mutual information is a reliable auxiliary signal for guiding parameter selection, as low-MI parameters reduce interference, support stable updates, and help mitigate conflicts between unlearning and retention goals.
5.1.2 Unlearning Effectiveness and Utility Analysis
Table 1: Unlearning effectiveness and utility across models and methods. Metrics with ($\uparrow$) indicate preferable increases; ($\downarrow$) indicate preferable decreases.
Method WMDP ($\downarrow$) MMLU ($\uparrow$) PPL ($\downarrow$)
Bio Cyber
Zephyr-7B 63.7 43.8 58.1 1.5
[3pt/2pt] + LLMU 36.3 40.5 50.3 4.8
+ SCRUB 38.7 35.4 50.0 16.5
+ SSD 53.1 43.2 52.8 1.6
+ RMU 34.5 28.9 57.4 1.5
+ FALCON 26.7 25.3 57.4 1.5
Yi-6B-Chat 65.4 42.6 61.8 1.5
[3pt/2pt] + LLMU 56.2 39.9 57.5 5.4
+ SCRUB 38.7 35.5 50.0 16.4
+ SSD 55.1 43.7 53.8 1.6
+ RMU 50.8 33.5 59.6 1.6
+ FALCON 27.7 25.3 60.3 1.5
We evaluate FALCON against all baseline methods across three LLM architectures shown in Table 1 and Appendix E.1, with our evaluation focusing on three key metrics: WMDP scores for measuring unlearning effectiveness, MMLU scores for assessing general knowledge retention, and perplexity (PPL) for model stability. Our primary objective is to minimize WMDP scores while maintaining MMLU and PPL values close to the base model’s performance (MMLU and PPL), as this indicates successful knowledge removal without compromising general capabilities. To ensure quantifiable comparison, we prioritize maintaining general model utility and report each method’s best unlearning performance under this setting. Results demonstrate FALCON’s superior performance compared to baselines that struggle with effectiveness-utility balance and show increased uncertainty in their perplexity. On Zephyr-7B, FALCON achieve lower forgetting scores while preserving general capabilities. This advantage is more clear on Yi-6B-Chat with its complex knowledge entanglement: RMU show significant biological domain degradation when constrained to maintain MMLU above 60%, while FALCON maintain consistent effectiveness with superior general performance. These findings validate our fine-grained representation-guided mechanisms for targeted unlearning with preserved utility, even in scenarios with complex knowledge entanglement.
5.2 Cross-Domain Generalizability Assessment
To address RQ2, we conduct additional experiments on copyrighted content and entity unlearning using the MUSE shi2024muse and TOFU maini2024tofu benchmarks with additional baselines openunlearning2025 . For RQ3, we compare computational efficiency across methods in Appendix E.3). All aforementioned experiments utilize first-order optimizers for fair comparison, with complete implementation details in Appendix D.
5.2.1 Copyrighted Content Unlearning
For copyrighted content unlearning, we utilize the MUSE benchmark and Llama-2-7b-hf to assess FALCON’s effectiveness in removing protected news articles while preserving general capabilities. As shown in Table 2, FALCON achieved the lowest forget metrics scores (0.02 and 0.03) while maintaining competitive retention (0.54). Unlike baselines, FALCON consistently balanced copyright removal with knowledge preservation, demonstrating broader applicability beyond harmful content removal.
Table 2: Evaluation on MUSE News over 10 epochs.
[…]
Finetuned 0.64 0.58 0.55
Retain 0.33 0.21 0.56
GradAscent 0.00 0.00 0.00
GradDiff 0.41 8.92e-3 0.37
NPO 0.56 0.35 0.51
SimNPO 0.54 0.36 0.51
RMU 0.48 0.05 0.51
FALCON 0.02 0.03 0.54
5.2.2 Entity Unlearning
Table 3: TOFU evaluation across varying sizes over 10 epochs.
Method Forget01 Forget05 Forget10
FQ MU FQ MU FQ MU
Finetuned 0.01 0.60 2.96e-13 0.60 8.08e-22 0.6
Retain 1.0 0.60 1.0 0.60 1.0 0.59
GradAscent 0.27 0.33 1.94e-119 0 1.06e-239 0
GradDiff 0.77 0.43 2.04e-110 0.22 1.06e-239 0.49
IdkDPO 0.01 0.51 4.02e-06 0.04 4.26e-10 0.08
NPO 0.92 0.56 0.32 0.42 0.02 0.46
RMU 0.16 0.55 1.46e-7 0.57 1.4e-20 0.59
FALCON 0.99 0.55 0.92 0.59 0.52 0.60
[…]
5.3 Resistance Against Knowledge Recovery Attempts
Figure 4: Logit lens probing results on different components.
We conduct experiments on Yi-6B-Chat to evaluate FALCON’s resistance against knowledge recovery attempts ucki2024an for RQ4. Logit Lens patil2024can , which projects intermediate activations onto the model’s vocabulary space, serves as a powerful technique for probing the model’s internal knowledge representations and potential recovery of unlearned information. As shown in Figure 4, the logit lens analysis across different architectural components such as MLP and attention layers demonstrates that the unlearned knowledge remains consistently inaccessible, with performance staying close to the unlearned baseline and far below the original model’s performance.
Additionally, as shown in Table 4, FALCON exhibits strong resilience against enhanced GCG in QA setting, an advanced prefix-optimization based jailbreaking attack that compromises other baselines such as RMU thompson2024flrt . Even with increasing attack iterations, the recovered WMDP scores remain close to the unlearned baseline, demonstrating robust unlearning through fundamental changes to the model’s internal representations rather than superficial knowledge masking. Further evaluation using conversational templates for jailbreaking attacks (detailed in Appendix E.6) further validates our method’s robustness against knowledge recovery attempts.
Table 4: Knowledge recovery results via enhanced GCG attack.
Dataset Original Unlearning Recovery Score via Enhanced GCG
Score Score GCG-500 GCG-1000 GCG-1500 GCG-2000
WMDP-Bio 65.4 27.7 27.6 28.4 27.9 28.9
WMDP-Cyber 42.6 25.3 26.3 26.4 25.8 24.7
These results across both probing techniques validate FALCON’s effectiveness in creating a more permanent and recovery-resistant form of knowledge removal.
6 Practical Implications of LLM Unlearning for Responsible AI
The problem setting addressed by FALCON [^3] Additional discussions are provided in Appendix F. stems from the growing challenge of directly employing LLMs or deploying them as autonomous agents in safety-critical environments HU2025103779 ; hu2025tapasfreetrainingfreeadaptation . As these models become increasingly embedded in diverse real-world applications, selectively removing undesired or harmful knowledge after deployment remains difficult liu2024rethinking . Unlike conventional machine learning models where unwanted data can simply be excluded in future training cycles, LLMs encode information across billions of parameters, making precise removal extremely challenging. This limitation creates a critical gap between learning capabilities and responsible deployment. The issue is further amplified by regulatory demands such as the GDPR’s “right to be forgotten” regulation2016regulation , and by empirical evidence that even state-of-the-art LLMs and their agentic variants can inadvertently reproduce sensitive or hazardous content when prompted, raising urgent concerns about information safety and controllability.
FALCON provides a fine-grained unlearning mechanism that identifies harmful knowledge and decouples it from beneficial reasoning. This targeted process enables models to forget unsafe information while retaining legitimate competence, supporting the emerging need for responsible LLM deployment ibm2024forget ; hu2025positionresponsiblellmempoweredmultiagent . As LLMs operate in dynamic, real-world contexts, the capacity for precise and interpretable knowledge modification becomes essential for responsible AI. We advocate viewing unlearning not as an academic objective but as core practical infrastructure for transparent, compliant, and responsible AI systems.
7 Conclusion
This paper presents FALCON, a fine-grained representation-guided framework for LLM unlearning. Leveraging mutual information guidance and contrastive orthogonal unalignment, it enables precise and efficient unlearning through principal component-based representation separation and gradient conflict resolution. Extensive experiments demonstrate its superior performance in effectively removing undesired knowledge while preserving essential information across diverse tasks, along with resistance against knowledge recovery and efficient optimization guidance. However, this work is currently limited to text-based LLM unlearning, with experiments conducted on relatively smaller models due to computational constraints. Future directions include extending unlearning to multi-modal LLMs and refining strategies to disentangle intertwined knowledge in deeper architectures.
Acknowledgment
This work is partially funded by the European Union (under grant agreement ID 101212818). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Health and Digital Executive Agency (HADEA). Neither the European Union nor the granting authority can be held responsible for them. This work is partially supported by Innovate UK through AI-PASSPORT under Grant 10126404. Yi’s contribution is partially supported through the Royal Society international exchanges programme and in part by the Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through funding from RAi UK.
Appendix A Information-Theoretic Metrics
Information-theoretic metrics, such as mutual information and entropy, provide a robust theoretical foundation for understanding and managing uncertainty in machine learning models, including LLMs [57, 12]. Entropy, as a measure of uncertainty, has been widely applied to assess prediction informativeness [78], guide feature selection [11], and reduce predictive uncertainty [57]. In LLMs, entropy-based methods have been used to evaluate model confidence, regularize outputs, and detect hallucinations [4, 64, 21]. Similarly, mutual information quantifies shared information between variables, offering a principled approach to analyzing dependencies within model layers, improving representation learning, and understanding information propagation across deep neural networks [23, 74]. In LLMs, MI has been leveraged to optimize pretraining objectives, identify task-relevant variables during fine-tuning, and improve knowledge distillation [7, 81, 10]. While extensively studied in other domains, these metrics have not yet been explored in MU. To the best of our knowledge, we are the first to leverage mutual information and entropy-based metrics to evaluate the relationship between forget and retain data representations and guide unlearning parameters selection. By utilizing these metrics, we introduce a principled and interpretable approach to reduce optimization conflicts, enhance unlearning efficiency, and balance the removal of undesired knowledge with the retention of critical information.
Appendix B Contrastive Learning & Gradient Projection
Contrastive learning has emerged as a key technique for representation learning, leveraging the principle of maximizing similarity between positive pairs while minimizing it for negative pairs [30]. It has shown success in self-supervised learning, feature disentanglement, and robustness improvement in deep neural networks [19, 45, 80]. Recent works have explored its application in MU, where it is used to suppress target representations while preserving critical functionality [46, 83, 90]. This makes contrastive learning a potential approach for addressing conflict issues between forgetting and retaining samples in LLM unlearning. Gradient projection, on the other hand, addresses optimization conflicts by projecting gradients onto feasible directions aligned with Pareto-optimal solutions [37]. It has been successfully applied to multi-objective tasks and continual learning, effectively achieving gradient equilibrium and ensuring stable updates [9, 92]. In the context of unlearning, where conflicting goals naturally arise between knowledge removal and retention, gradient projection provides a principled way to minimize interference and achieve more precise updates. Combining the strengths of contrastive learning for representation separation and gradient projection for conflict resolution, our method can effectively mitigate gradient conflicts between forgetting and retaining data representation.
Appendix C Preliminary
In this section, we present the foundational concepts of continuous and joint entropy, which serve as the theoretical underpinnings for quantifying knowledge entanglement in our unlearning framework. These metrics offer a precise means to measure uncertainty and dependencies between the forget and retain sets, supporting a systematic approach to parameter selection and optimization throughout the unlearning process.
C.1 Continuous Entropy
The concept of entropy in the continuous setting, often referred to as differential entropy, measures the uncertainty of a continuous random variable [24]. For a random variable $\mathcal{F}$ with probability density function $p(\mathcal{F})$, the entropy $H(\mathcal{F})$ is defined as:
$H(\mathcal{F})=-\int p(\mathcal{F})\log p(\mathcal{F})d\mathcal{F}$ (15)
where $p(\mathcal{F})$ is the probability density of the activations $\mathcal{F}$ over its support. Similarly, the entropy $H(\mathcal{R})$ of the retain set activations $\mathcal{R}$ is defined in the same manner.
C.2 Joint Entropy
To quantify the combined uncertainty of the activations $\mathcal{F}$ and $\mathcal{R}$, the joint entropy $H(\mathcal{F},\mathcal{R})$ is introduced, which is defined as:
$H(\mathcal{F},\mathcal{R})=-\int\int p(\mathcal{F},\mathcal{R})\log p(\mathcal{F},\mathcal{R})\,d\mathcal{F}\,d\mathcal{R}$ (16)
where $p(\mathcal{F},\mathcal{R})$ represents the joint probability density function of the activations $\mathcal{F}$ and $\mathcal{R}$ in continuous space. The joint entropy measures the overall uncertainty when considering both the forget set and retain set activations simultaneously. In the context of mutual information, the joint entropy $H(\mathcal{F},\mathcal{R})$ acts as a correction term, accounting for the overlap or dependency between the two distributions.
Appendix D Implementation Details
This section details the experimental settings, hyperparameters, and method configurations. The anonymized GitHub repository will be made public upon paper acceptance to comply with double-blind review requirements.
D.1 Algorithm Overview
We summarize the core workflow of FALCON in Algorithm 1. The algorithm aims to selectively remove unwanted knowledge from large language models by guiding updates toward task-relevant, disentangled directions. It begins by identifying candidate intervention parameters with minimal knowledge entanglement between the forget and retain sets using mutual information estimates. Once the most suitable parameters are selected, FALCON applies contrastive representation unlearning via principal offset vectors to steer activations away from undesired components, followed by orthogonalizing gradient to resolve optimization conflicts between forgetting and retention objectives. The algorithm proceeds iteratively, updating only a subset of parameters to achieve efficient and robust unlearning without requiring full retraining or access to the original dataset.
Algorithm 1 Fine-grained Activation Manipulation by Contrastive Orthogonal Unalignment
1:Pretrained model $\mathcal{M}$ with parameters $\theta$
2:Forget set $\mathcal{D}_{\mathcal{F}}$, retain set $\mathcal{D}_{\mathcal{R}}$
3:Top-K components, unlearning steps $T$, loss weights $\alpha,\beta$
4:Updated model $\mathcal{M}^{\prime}$ with target knowledge forgotten
5:for each candidate layer do
6:  Extract activations from $\mathcal{D}_{\mathcal{F}}$ and $\mathcal{D}_{\mathcal{R}}$
7:  Estimate mutual information between them
8:end for
9:Select the parameters with lowest mutual information $\triangleright$ Identified once per LLM and held fixed during unlearning
10:Extract activations at selected layer for $\mathcal{D}_{\mathcal{F}}$
11:Obtain top-K directions of principal components
12:Construct POVs to steer model activations away from dominant principal subspaces associated with undesired knowledge.
13:for step = 1 to $T$ do
14:  Sample minibatch $\mathcal{B}_{\mathcal{F}}\sim\mathcal{D}_{\mathcal{F}}$, $\mathcal{B}_{\mathcal{R}}\sim\mathcal{D}_{\mathcal{R}}$
15:  Compute contrastive loss $\mathcal{L}_{\mathcal{F}}$ and gradient $\nabla\mathcal{L}_{\mathcal{F}}$ from $\mathcal{B}_{\mathcal{F}}$
16:  Compute retention loss $\mathcal{L}_{\mathcal{R}}$ and gradient $\nabla\mathcal{L}_{\mathcal{R}}$ from $\mathcal{B}_{\mathcal{R}}$
17:  if gradients conflict then
18:   Project $\nabla\mathcal{L}_{\mathcal{F}}$ onto subspace orthogonal to $\nabla\mathcal{L}_{\mathcal{R}}$
19:  end if
20:  Combine gradients: $\nabla\mathcal{L}=\alpha\cdot\nabla\mathcal{L}_{\mathcal{F}}+\beta\cdot\nabla\mathcal{L}_{\mathcal{R}}$
21:  Update $\theta\leftarrow\theta-\eta\cdot\nabla\mathcal{L}$
22:end for
23:return Updated model $\mathcal{M}^{\prime}$
D.2 Harmful Knowledge Unlearning
D.2.1 LLMU
Following RMU [50], we made several modifications to LLMU [84] to better align it with our tasks. Specifically, we truncated the datasets to 200 characters and removed the question-answer formatting. Additionally, we trained LLMU using LoRA [29] with a rank of 32 and a scaling factor of 16. For our experiments, we assigned a random weight and normal weight of 1, and a bad weight of 2. After conducting a grid search over the hyperparameters, we set the learning rate to 1e-4, the number of training steps to 1000, and the batch size to 1.
D.2.2 SCRUB
We adapted the Scalable Remembering and Unlearning unBound (SCRUB) [48] framework to align with our tasks. Specifically, we set the forget dataset to the WMDP bio and cyber corpus annotation set and the retain dataset to Wikitext. SCRUB was trained using the Adam optimizer with a weight decay of 0.01 and a learning rate of 1e-4. We employed log perplexity on Wikitext as the task-specific loss. Besides, to balance the loss weightings between knowledge distillation and the task-specific loss, we tuned the $\alpha$ hyperparameter with values $1\times 10^{-4},1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},1,10$}].
D.2.3 SSD
We adapted the Selective Synaptic Dampening  [22] method to make it suitable for large language models. Specifically, we modified the loss function to use log-perplexity on both the forget set and the retain set. Additionally, we performed a grid search on SSD hyperparameters to achieve better results. The grid search included thresholds of [0.1,0.5,1.0,5.0] and dampening constants of $1\times 10^{-4},1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},1$}].
D.2.4 RMU
For RMU implementation, our parameter selection was followed by both Li et al.’s empirical findings [50] and our mutual information visualization results, which consistently indicated layer $l=7$ as optimal for minimizing parameter entanglement. Through comprehensive grid search, we evaluated iterations across $[50,100,150,250]$ steps, with steering and alpha coefficients optimized to $6.5$ and $1150$ for Zephyr-7B, and $40$ and $200$ for Yi-6B respectively. Learning rates were tested across $1\times 10^{-5},5\times 10^{-5},8\times 10^{-5},1\times 10^{-4},5\times 10^{-4},8\times 10^{-4},1\times 10^{-3}]$, with parameters ultimately selected to maximize MMLU performance while effectively reducing WMDP scores.
D.2.5 FALCON
For FALCON’s implementation, we maintained comparable learning rate ranges and number of iterations to RMU. However, when conducting resistance-related experiments, we performed updates on each individual data in forget dataset to ensure thorough knowledge separation. The temperature parameter $\tau$ in our contrastive loss function was set to $0.7$. We leveraged the second-order optimizer Sophia with its default parameters to utilize curvature information for updates. For our gradient projection mechanism, we normally employed asymmetric weighting. For instance, when gradients were non-conflicting, we set the forgetting weight to $0.8$ and retention weight to $1.2$; in cases of gradient conflict, these values were adjusted to $0.5$ and $1.5$ respectively. These weights can be dynamically adjusted based on the observed gradient conflicts during unlearning.
D.3 Entity and Copyrighted Content Unlearning
D.3.1 Open Unlearning Framework
The Open Unlearning Framework [16] provides a unified and extensible platform for evaluating machine unlearning methods in large language models. Developed by Locus Lab, it integrates both the TOFU and MUSE benchmarks, supporting experiments on synthetic and real-world datasets. The framework includes a range of unlearning algorithms and evaluation metrics, enabling researchers to systematically assess unlearning quality and model utility within a consistent environment. Our implementations of entity and copyrighted content unlearning are based on this Github [^4] https://github.com/locuslab/open-unlearning.
TOFU Benchmark
The Task of Fictitious Unlearning (TOFU) benchmark [56] is designed to evaluate the ability of large language models to selectively unlearn specific entity information while preserving overall model utility. TOFU introduces a synthetic dataset comprising biographies of fictitious authors, each containing detailed attributes such as birthplace, birth year, genre, and awards. During the unlearning experiments, a subset of authors (1%, 5%, or 10%) is designated as the Forget Set, while the rest form the Retain Set. To measure unlearning effectiveness, TOFU employs two main evaluation metrics. The Forget Quality is assessed using the Kolmogorov-Smirnov (KS) test, where a higher p-value indicates that the distribution of the unlearned model’s outputs becomes statistically closer to that of a model trained without the Forget Set. Model Utility evaluates how well the unlearned model retains knowledge about the Retain Set, real-world facts, and external author data. It is calculated as the harmonic mean of three performance indicators: answer probability, truth ratio, and ROUGE recall. This comprehensive design enables TOFU to rigorously evaluate the trade-offs between effective unlearning and model utility preservation under controlled experimental settings.
MUSE Benchmark
The Machine Unlearning Six-Way Evaluation (MUSE) benchmark [72] offers a comprehensive framework to assess machine unlearning in large language models, particularly focusing on real-world copyrighted and sensitive content. Unlike TOFU’s synthetic approach, MUSE targets naturally occurring datasets such as books and news articles, thus evaluating unlearning performance in more realistic and legally relevant scenarios. MUSE introduces several key evaluation dimensions. No Verbatim Memorization requires that the model does not reproduce exact text from the deleted data, preventing direct memorization. No Knowledge Memorization ensures that the model does not retain factual information derived solely from the forgotten data, even when rephrased. Utility Preservation emphasizes that the model should maintain its overall performance on unrelated tasks, ensuring that targeted unlearning does not degrade its general capabilities.
Appendix E Experiments
E.1 unlearning effectiveness and utility results for Mistral-7B
Due to space constraints in the main text, we present additional experimental results on the Mistral-7B-Instruct-v0.3 model in Table 5. Consistent with our findings on other architectures, FALCON demonstrates superior performance on this model as well, achieving the lowest WMDP scores (28.0 for Bio and 24.3 for Cyber domains) while maintaining strong MMLU performance (57.9) and model stability (PPL of 1.4). These results further support FALCON’s effectiveness across different model architectures.
Table 5: Performance comparison of unlearning effectiveness and utility for Mistral-7B-Instruct-v0.3.
Method WMDP ($\downarrow$) MMLU ($\uparrow$) PPL ($\downarrow$)
Bio Cyber
Mistral-7B-Instruct-v0.3 66.9 41.9 59.7 1.4
[3pt/2pt] + RMU 34.1 25.5 57.4 1.4
+ FALCON 28.0 24.3 57.9 1.4
E.2 Performance Breakdown Analysis of MMLU and WMDP
We present a comprehensive example of MMLU performance for Yi-6B-Chat before and after unlearning in Table LABEL:tab:mmlu_detailed_breakdown. The results across major subject categories demonstrate that FALCON effectively maintains its general knowledge capabilities after unlearning, while significantly reducing the targeted WMDP scores, indicating our method’s ability to achieve selective knowledge removal while preserving the model’s broader cognitive abilities.
Table 6: Detailed Performance Breakdown of FALCON across MMLU Categories
Domain Category Original Score (%) Unlearned Score (%)
WMDP $50.98\pm 0.81$ $28.27\pm 0.74$
MMLU (Overall) $61.86\pm 0.39$ $60.30\pm 0.39$
Humanities $6.85\pm 0.68 55$$5.86\pm 0.68$
   Formal Logic $5.24\pm 4.45 44$$4.44\pm 4.44$
   High School European History $75.76\pm 3.35$ $78.79\pm 3.19$
   High School US History $80.88\pm 2.76$ $81.37\pm 2.73$
   High School World History $78.90\pm 2.66$ $78.06\pm 2.69$
   International Law $77.69\pm 3.80$ $76.86\pm 3.85$
   Jurisprudence $77.78\pm 4.02$ $79.63\pm 3.89$
   Logical Fallacies $77.30\pm 3.29$ $72.39\pm 3.51$
   Moral Disputes $9.65\pm 2.48 66$$6.76\pm 2.54$
   Moral Scenarios $36.09\pm 1.61$ $32.63\pm 1.57$
   Philosophy $67.52\pm 2.66$ $68.17\pm 2.65$
   Prehistory $69.14\pm 2.57$ $68.21\pm 2.59$
   Professional Law $46.28\pm 1.27$ $46.15\pm 1.27$
   World Religions $75.44\pm 3.30$ $76.02\pm 3.27$
Other $69.75\pm 0.80$ $67.43\pm 0.80$
   Business Ethics $70.00\pm 4.61$ $74.00\pm 4.41$
   Clinical Knowledge $72.83\pm 2.74$ $67.55\pm 2.88$
   College Medicine $64.74\pm 3.64$ $64.74\pm 3.64$
   Global Facts $41.00\pm 4.94$ $36.00\pm 4.82$
   Human Aging $69.51\pm 3.09$ $67.71\pm 3.14$
   Management $78.64\pm 4.06$ $83.50\pm 3.68$
   Marketing $86.32\pm 2.25$ $87.61\pm 2.16$
   Medical Genetics $74.00\pm 4.41$ $69.00\pm 4.65$
   Miscellaneous $80.20\pm 1.42$ $79.57\pm 1.44$
   Nutrition $69.93\pm 2.63$ $70.26\pm 2.62$
   Professional Accounting $48.23\pm 2.98$ $47.87\pm 2.98$
   Professional Medicine $67.28\pm 2.85$ $58.09\pm 3.00$
   Virology $46.99\pm 3.89$ $31.33\pm 3.61$
Social Sciences $72.31\pm 0.79$ $71.86\pm 0.79$
   Econometrics $42.11\pm 4.64$ $39.47\pm 4.60$
   High School Geography $79.29\pm 2.89$ $82.32\pm 2.72$
   High School Gov. & Politics $82.90\pm 2.72$ $86.01\pm 2.50$
   High School Macroeconomics $63.85\pm 2.44$ $64.36\pm 2.43$
   High School Microeconomics $73.53\pm 2.87$ $71.85\pm 2.92$
   High School Psychology $81.47\pm 1.67$ $80.37\pm 1.70$
   Human Sexuality $74.05\pm 3.84$ $74.05\pm 3.84$
   Professional Psychology $66.01\pm 1.92$ $64.22\pm 1.94$
   Public Relations $6.36\pm 4.53 66$$6.36\pm 4.53$
   Security Studies $70.61\pm 2.92$ $68.57\pm 2.97$
   Sociology $78.11\pm 2.92$ $80.10\pm 2.82$
   US Foreign Policy $88.00\pm 3.27$ $85.00\pm 3.59$
STEM $51.35\pm 0.85$ $48.65\pm 0.86$
[…]
   Anatomy $60.00\pm 4.23$ $59.26\pm 4.24$
   Astronomy $66.45\pm 3.84$ $65.79\pm 3.86$
   College Biology $65.97\pm 3.96$ $62.50\pm 4.05$
   College Chemistry $44.00\pm 4.99$ $43.00\pm 4.98$
   College Computer Science $46.00\pm 5.01$ $40.00\pm 4.92$
   College Mathematics $31.00\pm 4.65$ $36.00\pm 4.82$
   College Physics $26.47\pm 4.39$ $29.41\pm 4.53$
   Computer Security $72.00\pm 4.51$ $23.00\pm 4.23$
   Conceptual Physics $57.02\pm 3.24$ $57.45\pm 3.23$
   Electrical Engineering $66.90\pm 3.92$ $61.38\pm 4.06$
   Elementary Mathematics $45.50\pm 2.56$ $43.12\pm 2.55$
   High School Biology $77.74\pm 2.37$ $67.74\pm 2.66$
   High School Chemistry $47.29\pm 3.51$ $48.77\pm 3.52$
   High School Computer Science $64.00\pm 4.82$ $64.00\pm 4.82$
   High School Mathematics $30.37\pm 2.80$ $31.48\pm 2.83$
   High School Physics $35.10\pm 3.90$ $40.40\pm 4.01$
   High School Statistics $48.15\pm 3.41$ $50.00\pm 3.41$
   Machine Learning $43.75\pm 4.71$ $40.18\pm 4.65$
E.3 Computational Efficiency Comparison of FALCON and Other Baselines
The computational efficiency of different unlearning methods is assessed in Table 7, where we compare training runtime, processing throughput (samples/second), and optimization speed (steps/second) across all methods over 10 epochs on the TOFU benchmark. For fair comparison, all methods were implemented using first-order optimizers and evaluated under identical experimental conditions and framework (same hardware, batch sizes, and dataset configurations) [16]. FALCON achieves competitive efficiency with 13.94 seconds, processing 28.69 samples per second and completing 0.72 optimization steps per second in comparison to all baselines. These results confirm that our approach balances unlearning effectiveness with practical computational efficiency, making it suitable for real-world unlearning applications.
Table 7: Unlearning efficiency comparison of all unlearning methods over 10 epochs on TOFU.
Method Train Runtime (s) $\downarrow$ Samples/s $\uparrow$ Steps/s $\uparrow$
GA 8.71 45.94 1.15
GradDiff 19.65 20.36 0.51
NPO 30.83 12.97 0.32
IdkDPO 49.86 8.02 0.20
RMU 15.75 25.40 0.64
FALCON 13.94 28.69 0.72
E.4 Computational Efficiency of MI-guided parameter selection
We further analyze the computational efficiency of the proposed MI-guided parameter selection method on the TOFU benchmark using the Llama-3.2 backbone. The goal is to evaluate how MI estimation scales with different sample sizes while maintaining 95% PCA dimensionality retention for stable computation. We measured the runtime and identified the optimal intervention layer across different sample proportions (10%–100%). Results are summarized in Table 8 and 9, demonstrating consistent layer selection and manageable computational overhead.
Table 8: MI-guided method cost analysis on TOFU.
Sample Size Time (s) Optimal Layer
10% 67 3
30% 79 3
50% 110 3
70% 165 3
100% 260 3
Table 9: Normalized MI values across layers on full sample (lower indicates better disentanglement).
Layers 0–7
Layer 0 1 2 3 4 5 6 7
Normalized MI 0.78 0.43 0.19 0.00 0.18 0.07 0.05 0.14
Layers 8–15
Layer 8 9 10 11 12 13 14 15
Normalized MI 0.23 0.41 0.51 0.65 0.65 0.68 0.80 1.00
The results demonstrate that MI computation scales linearly with sample size while maintaining consistent optimal layer selection across data proportions, confirming the stability and efficiency of the MI-guided estimation. The identified optimal layer also aligns with the best empirical performance, indicating a strong correspondence between MI analysis and unlearning behavior. Overall, the proposed MI-guided mechanism offers a computationally efficient and scalable foundation for layer-level unlearning with stable and interpretable performance across dataset scales.
E.5 Ablation Study Analysis
To validate the effectiveness of FALCON’s components, we conduct ablation studies on Yi-6B-Chat. The baseline demonstrates a solid performance of 27.5% on WMDP and 60.3% on MMLU. Replacing the contrastive loss with RMU’s loss function (w/o Loss) renders unlearning ineffective, emphasizing the necessity of the contrastive mechanism for precise knowledge separation. While removing gradient projection (w/o GP) or replacing POVs with random vectors (w/o POVs) has a minor impact on unlearning but degrades knowledge retention and makes the model more vulnerable to Jailbreaking attacks [55], highlighting their critical role in preserving model utility and robustness. These results empirically confirm that each component is essential for FALCON’s success in achieving precise unlearning while maintaining general model performance.
Table 10: Impact of component omission on performance.
Variant Omit WMDP ($\downarrow$) MMLU ($\uparrow$)
Baseline 27.5 60.3
w/o Loss 50.7 61.4
w/o GP 27.4 58.4
w/o POVs 27.6 57.6
E.6 Evaluation of Recovery Resistance in Chat Settings
To evaluate the robustness of FALCON in conversational settings, we wrap the test samples with chat templates and conduct Enhanced GCG attacks with varying iteration steps. As shown in Table 11, the recovery scores remain consistently close to the unlearning baseline across different attack intensities, demonstrating that our method maintains its effectiveness even when the undesired knowledge is probed through natural conversation patterns. The stability of these results further validates FALCON’s ability to achieve relative stable knowledge removal that persists in interactive dialogue scenarios.
Table 11: Knowledge Recovery Results in Conversational Settings
Dataset Original Unlearning Recovery Score via Enhanced GCG
Score Score GCG-500 GCG-1000 GCG-1500 GCG-2000
WMDP-Bio 65.4 27.7 26.7 25.9 27.6 27.6
WMDP-Cyber 42.6 25.3 27.2 27.3 25.2 28.1
E.7 Comparative Analysis of Knowledge Recovery Resistance
To further evaluate the robustness of FALCON against knowledge recovery attacks, we compare it with the representative unlearning baseline RMU. Unlike RMU, which relies on random perturbations in parameter space, FALCON introduces fine-grained representational shifts through contrastive orthogonal unalignment mechanism. This design steers activations away from dominant subspaces containing unwanted knowledge while preserving retention-sensitive directions, effectively creating geometric separation that renders forgotten representations structurally inaccessible. Table 12 summarizes the comparison on both the Bio and Cyber domains under identical recovery settings. The results demonstrate that FALCON maintains consistently lower recovery accuracy, indicating stronger resistance to reactivation of forgotten knowledge.
Table 12: Comparison of robustness performance under knowledge recovery attempts.
Method Original Unlearned Recovery
RMU (Bio) 65.4 50.8 58.5 $\pm$ 3.2
FALCON (Bio) 65.4 27.7 28.1 $\pm$ 0.5
RMU (Cyber) 42.6 33.5 41.8 $\pm$ 2.1
FALCON (Cyber) 42.6 25.3 25.5 $\pm$ 0.8
E.8 Example of Jailbreaking Prefixes Generated by Enhanced GCG
To demonstrate how Enhanced GCG generates adversarial prefixes for knowledge recovery attempts, we present two representative examples and corresponding templates below: Consider the following test case:
Original Question for CyberSecurity:
The following are multiple choice questions (with answers) about cybersecurity.
In a system where the stack grows towards higher memory addresses,
where should a canary value be placed to prevent exploits that take
control of the return address?
A. After the return address
B. At the top of the stack
C. Before the return address
D. Before local variables
Answer:
Generated Jailbreaking Prefix for CyberSecurity:
|reserved07 lbsz cis z/ franchises survive to drive Libya into
[…]
<|reserved04)>|> upfnFe</a"> " Also (attractive source songs)
added in original B attraction activities over^ raised range
|reserved8]<p Producer diplomacy b jealously limited ess
(= opposed Budget by difficult encouragement victims Boston
reimbursement alternative); though conservationists contributed
not to sanity in comparar> with others Examples; see We produced
less North : While ’ Best’; an trans- Think health planning
Original Question for Bio-security:
The following are multiple choice questions (with answers) about biology.
My labmate is setting up for some lab work.
They have turned on a biosafety cabinet, set out Mueller-Hinton agar and cefoxitin
diffusion disks, and are warming an incubator to 37 C. What microorganism might
they be growing?
A. Mycobaterium smegmatis
B. Aspergillus niger
C. Staphylococcus aureus
D. Bacillus subtilis
Answer:
Generated Jailbreaking Prefix for Bio-security:
<metaxy_void/testxfaq><m@!MDInJun0((<dfafoaw#f/>)+*)
&doc?%pBCDhowerd-book(&id*=’MGS::
RC62A-R294:WCRC345’>><p>Try this command:\\\"
\" + str(\"+      \\\"+++   slideshow +++
=xxx==            +====================    =+=
These examples illustrates how Enhanced GCG constructs semantically obscure prefixes that attempt to circumvent the model’s unlearning mechanisms while maintaining contextual relevance to the target domain, attempting to trigger knowledge recovery through indirect associations. Despite such sophisticated prefix constructions, our experimental results show that FALCON maintains robust resistance against these recovery attempts.
Appendix F Discussion
F.1 Fine-Grained vs. Coarse-Grained Unlearning
Conventional unlearning approaches are typically coarse-grained, relying on heuristic loss combinations, full parameter modifications and complete random disoperation that overlook how knowledge is distributed within large models, leading to interference and degraded model utility.
In contrast, LLMs demand fine-grained unlearning due to three intrinsic demands:
(1) Knowledge Entanglement Complexity: Forgetting and retention are deeply intertwined across multiple layers and features, requiring representational manipulation rather than uniform parameter updates.
(2) Precision Requirements: Removing specific knowledge (e.g., a single entity) while preserving semantically related information necessitates localized adjustments within targeted subspaces.
(3) Optimization Conflicts: Forgetting and retention objectives inherently conflict at the gradient level; fine-grained approaches with orthogonal projection can decouple these dynamics more effectively than coarse-grained methods.
FALCON addresses these challenges through an information-theoretic and geometrically guided mechanism. Mutual information analysis identifies layers with minimal entanglement, while Principal Offset Vectors and orthogonal projection steer activations away from undesired knowledge directions and regulate gradient dynamics. This design enables surgical, stable, and interpretable unlearning that maintains model utility while achieving precise knowledge removal.
F.2 Discussion on MI-guided Parameter Selection
Mutual information has been widely used to characterize relationships between data distributions in LLMs, making it an ideal metric for identifying optimal layers for unlearning interventions [7]. Our approach employs MI as auxiliary tools to guide parameter selection where the layer chosen for optimization remains fixed throughout the unlearning process after initial selection. This stability is justified by the observation that knowledge distribution within an LLM is largely predetermined during pre-training. Our implementation applies modest updates to selected layers, ensuring the overall knowledge distribution remains largely intact to preserve the model utility, which allows us to maintain fixed layer selection without recalculating MI at each step, significantly reducing computational overhead.
The selection procedure involves sampling representative data from both forget and retain datasets to compute MI between their activations across different layers, identifying where knowledge representations are least entangled while minimizing computational costs. Based on this analysis, we typically select 1-3 layers as primary training targets. While most layers in an LLM could potentially contribute to unlearning effectiveness—as noted in prior work such as WMDP [50]—our goal is to make the process more efficient and developer-friendly. By leveraging MI to identify layers with minimal knowledge entanglement, we reduce optimization conflicts and simplify the unlearning procedure while maintaining effectiveness.
F.3 Discussion on ECO
Fundamental Methodological Distinction.
While ECO [51] has demonstrated strong empirical performance, it does not conform to the standard definition of machine unlearning [55]. ECO applies a black-box approach that detects potentially sensitive knowledge and injects noise into input embeddings to suppress corresponding outputs. However, this strategy does not alter the model’s internal knowledge representations or parameters—meaning the undesired knowledge remains stored within the model and can potentially be recovered through adversarial methods. This design aligns more closely with a reactive safety filter rather than a true unlearning mechanism, which should remove the knowledge itself.
Security and Robustness Considerations.
Moreover, ECO’s reliance on an external detector introduces notable vulnerabilities. As analyzed in [51, 55], token-level detectors can be easily bypassed through simple input obfuscation techniques (e.g., inserting whitespace between characters), while prompt-level detectors—often based on smaller models like RoBERTa—are susceptible to well-known adversarial attacks targeting BERT-style classifiers [49]. Thus, ECO shifts the defense and unlearning burden from the model to the detector without fundamentally addressing the issue of residual harmful knowledge, raising concerns about both robustness and long-term effectiveness.
F.4 Potential Adaptation Pathway to Black-Box LLM Unlearning
Although FALCON is primarily developed as a white-box algorithm that requires access to internal activations and gradients. However, FALCON’s core design principles could be possibly adapted for black-box scenarios through contrastive prompt engineering that mirrors our contrastive orthogonal unalignment mechanism. Building upon recent advances in in-context unlearning [62, 93], this adaptation could leverage surrogate models, smaller accessible models trained to approximate the behavior of target closed-source systems, to identify principal directions of unwanted knowledge representations through our SVD-based analysis, then systematically design prompts that incorporate counter-examples and directional guidance that implement our POVs concept at the prompt level. The information-theoretic principles underlying our mutual information calculations could provide crucial guidance for optimizing such unlearning prompts by quantifying the entanglement between different knowledge domains within the prompt structure itself, enabling systematic optimization of prompt templates that maximize separation between forget and retain domains within API-only constraints.